### v1

# Towards Black-Box Membership Inference Attack for Diffusion Models

Abstract
Identifying whether an artwork was used to train a diffusion model is an important research topic, given the rising popularity of AI-generated art and the associated copyright concerns. The work approaches this problem from the membership inference attack (MIA) perspective. We first identify the limitations of applying existing MIA methods for copyright protection: the required access of internal U-nets and the choice of non-member datasets for evaluation. To address the above problems, we introduce a novel black-box membership inference attack method that operates without needing access to the model’s internal U-net. We then construct a DALL-E generated dataset for a more comprehensive evaluation. We validate our method across various setups, and our experimental results outperform previous works.
1 Introduction
Recently, there has been a surge in the popularity of generative models, with diffusion models in particular, gaining huge attention within the AI community [38, 41, 42]. These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation [12, 40], text-to-image generation [32, 49, 29] and image-to-image generation [33]. This surge has given rise to powerful AI art models such as DALL-E 2 [31], Stable Diffusion [32] and Imagen [34]. The field of AI-generated art holds a promising future and is expected to have widespread impact.
Effective training of diffusion models requires high-quality data. It is thus crucial to design an algorithm that can identify whether a specific artwork has been used during the training of a model, thereby providing protection for these artworks and detecting misuse of data. This is a classic problem in the field of machine learning, first introduced by [37] and named “membership inference attack”.
A series of studies have been conducted on membership inference attacks against diffusion models. [14] was the first to examine this issue, utilizing the loss function values of diffusion models to determine whether an image is in the training set. [5] and [18] extended this work by relaxing the assumptions about model access requirements.
Despite great progress, previous methods are not yet ready for copyright protection in diffusion models for two reasons. First, most existing approaches heavily rely on checking whether the U-net of the model predicts noise accurately, which is not practical since most commercial diffusion models available today offer only API access, while the U-net remains hidden. We refer to MIA with access only to generated images as the Black-Box MIA. Second, existing evaluations for stable diffusion models primarily focus on content shifts in the data, while the shifts in art styles are not covered (illustrated in Figure 3).
To address the above issues, we propose a black-box membership inference attack method that does not require access to the U-net of the model. We posit that if we alter an image using the target diffusion model’s variation API, the sampling process will be captured by a region of “attraction” if the model has seen this image during training (illustrated in Figure 2). By quantifying the disparity between the resulting reconstructed image with the original, we can effectively determine whether the image was part of the training dataset.
We also construct a new evaluation dataset named LAION by DALL-E reflecting style shifts in data rather than content shifts. In addition to evaluating our methods on existing datasets across multiple models, we test MIA on the stable diffusion model to classify LAION by DALL-E from the ground truth training data LAION-5B.
The main contributions of this work are listed as follows:
  1. 1.
    We propose a black-box membership inference attack method that does not require access to the model’s internal structure. Instead, we leverage the unique properties of diffusion models. Our method only involves using the model’s variation API to alter an image and compare it with the original one. We name our method ReDiffuse.
  2. 2.
    We evaluate our method using DDIM and Stable Diffusion models on classical datasets, including CIFAR10, STL10-Unlabeled [2], LAION-5B [36], etc. Our method is shown to be effective and robust across datasets, and it outperforms the previous methods.
  3. 3.
    We propose a new MIA evaluation dataset which is complementary to previous work. We use the image titles from LAION-5B as prompts for DALL-E’s API [31] to generate images of the same contents but different styles. Under this setup, previous membership inference attack algorithms are no longer effective, while our method remains robust.
2 Related Works
Diffusion Model
The diffusion model, initially proposed by [38], has achieved remarkable results in producing high-quality samples across a variety of domains. This ranges from image generation [41, 42, 4], audio synthesis [30, 19, 15], and video generation [11, 13, 46], etc. Among existing diffusion models, the Denoising Diffusion Probabilistic Model (DDPM) [12] is one of the most frequently adopted. This approach introduces a dual-phase process for image generation: initially, a forward process gradually transforms training data into pure noise, followed by a reverse process that meticulously reconstructs the original data from this noise. Building on this model, there have been numerous follow-up studies, such as Stable Diffusion [32], which compresses images into a latent space and generates images based on text, and the Denoising Diffusion Implicit Models (DDIM) [40], which removes Gaussian randomness to accelerate the sampling generation process. These advancements demonstrate the versatility and potential of diffusion models.
Data Safety and Membership Inference Attack
In the era of big data, preserving data privacy is paramount, especially in training diffusion models using sensitive datasets like artists’ artworks, which are protected by copyright laws. Membership inference attacks, initially introduced by [37], serve as an effective means to detect potential misuse of data without proper authorization. Its objective is to ascertain whether a particular data sample participated in the training phase of a target model. This approach is instrumental in probing privacy breaches and identifying illicit data utilization. Researchers primarily focus on membership inference attacks for classification models [35, 48, 24, 22], embedding models [39, 6, 26], and generative models [8, 10, 1].
In the domain of membership inference attacks against diffusion models, [47, 14] use a white-box approach, which assumes access to the entire diffusion model and utilizes loss and likelihood to determine whether a sample is in the training set. More recently, [5, 18, 43] have relaxed these requirements, eliminating the need for the entire model. They leverage the insight that samples within the training set yield more accurate noise predictions, thereby achieving high accuracy in membership inference attack tasks. However, they also require the outputs of the U-net, as it is necessary to obtain the noise predictions of intermediate steps.
3 Preliminary
In this section, we begin by discussing several popular diffusion models. We first introduce the Denoising Diffusion Probabilistic Model (DDPM) [12]. Then, we extend to the Denoising Diffusion Implicit Model (DDIM) [40], which is a variant of DDPM used to accelerate image generation. Lastly, we discuss the Stable Diffusion [32], a model that condenses images into a latent space and generates images grounded in text descriptions.
Denoising Diffusion Probabilistic Model (DDPM)
A diffusion model provides a stochastic path between an image and noise. The forward process (denoted as $q$) iteratively incorporates Gaussian noise into an image, while the reverse process (denoted as $p_{\theta}$) gradually reconstructs the image from noise.
$q(x_{t}\mid x_{t-1})$ $=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})\,,$
$p_{\theta}(x_{t-1}\mid x_{t})$ $=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))\,,$
where $\mu_{\theta}(\cdot)$ and $\Sigma_{\theta}(\cdot)$ are the mean and covariance of the denoised image parameterized by the model parameters $\theta$, and $\beta_{t}$ is a noise schedule that controls the amount of noise added at each step.
Denoising Diffusion Implicit Model (DDIM)
DDIM modifies the sampling process to improve efficiency while maintaining high-quality image generation. Unlike DDPM, which requires a large number of denoising steps, DDIM uses a non-Markovian process to accelerate sampling.
$x_{t-1}=\phi_{\theta}(x_{t},t)=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x_{t},t)\,,$ (1)
where $\bar{\alpha}_{t}=\prod_{k=0}^{t}\alpha_{k}$, $\alpha_{t}+\beta_{t}=1$ and $\epsilon_{\theta}(x_{t},t)$ is the noise predicted by the model at step $t$. This formulation requires fewer sampling steps without compromising the quality of the generated images.
Stable Diffusion
Stable Diffusion leverages a variational autoencoder (VAE) [17] to encode images into a latent space and perform diffusion in this compressed space. The model uses a text encoder to guide the diffusion process, enabling text-to-image generation:
$z_{t-1}\sim p_{\theta}(z_{t-1}\mid z_{t},\tau_{\theta}(y))\,,\quad x=\text{Decoder}(z_{0})\,,$
where $x$ represents the output image, $z_{t}$ represents the latent variable at step $t$, and the text conditioning $\tau_{\theta}(y)$ is incorporated into the denoising process to generate the image. This approach significantly reduces computational costs and allows for high-quality image synthesis from textual descriptions.
4 Algorithm Design
In this section, we introduce our algorithm. We begin by discussing the definition of black-box API and the limitations of previous membership inference attack methods. In our formulations we assume DDIM as our target model. The formulations for DDPM is highly similar and we omit it for brevity. We will discuss the generalization to the latent diffusion model in Section 5.
4.1 Towards Black-Box MIA for Diffusion Models
Most previous works on membership inference attacks against diffusion models aim to prevent data leakage, and hence can rely on thresholding the model’s training loss. For instance, [14] involves a direct comparison of image losses, while [5, 18] evaluates the accuracy of the model’s noise prediction at initial or intermediate steps. However, the required access to the model’s internal U-net structure prevent applications to copyright protection because most servers typically provide only black-box API access.
We denote a model API as black-box if it only requires the input image and returns the output image. Specifically, we do not have access to the internal structure of the model. Below we formulate the definition of the black-box variation API used in our algorithm.
Definition 1 (The black-box variation API).
We define the variation API $V_{\theta}(x,t)$ of a model as follows. Suppose we have an input image $x$, and the diffusion step of the API is $t$. The variation API randomly adds $t$-step Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to the image and denoises it using the DDIM sampling process $\phi_{\theta}(x_{t},t)$ as described in Equation 1, returning the reconstructed image $V_{\theta}(x,t)=\hat{x}$. The details are as follows:
$x_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\quad\hat{x}=\Phi_{\theta}(x_{t},0)=\phi_{\theta}(\cdots\phi_{\theta}(\phi_{\theta}(x_{t},t),t-1),0)\,.$ (2)
This definition aligns with the image-to-image generation method of the diffusion model, making access to this API practical in many setups [25, 33, 45]. Some APIs provide the user with a choice of $t$, while others do not and use the default parameter. We will discuss the influence of different diffusion steps in Section 6.4, showing that the attack performances are relatively stable and not sensitive to the selection of $t$. We also note that the for the target model, we can substitute $\phi_{\theta}(x_{t},0)$ with other sampling methods, such as the Euler-Maruyama Method [27] or Variational Diffusion Models [16].
Figure 1: The overview of ReDiffuse. We independently input the image to the black-box variation API $n$ times with diffusion step $t$. We take the average of the output images and compare them with the original ones. If the difference is below a certain threshold, we determine that the image is in the training set.
4.2 Algorithm
In this section, we present our algorithm for black-box MIA. Our key intuition comes from the reverse SDE dynamics in continuous diffusion models [41, 42]. Specifically, one simplified form of the reverse SDE (i.e., the denoise step) has the formulation below:
$\dot{X_{t}}=(X_{t}/2-\nabla_{x}\log p(X_{t}))+dW_{t}\,,\quad t\in[0,T]\,.$ (3)
where $W_{t}$ denotes a standard Brownian motion (with abuse of notation $t$) and $T$ is the generation time.
Figure 2: The intuition of our algorithm design. We denote $x$ as the target image, $x_{i}$ as the $i$-th image generated by the variation API, and $\hat{x}$ as the average image of them. For member image $x$, the difference $v=x-\hat{x}$ will be smaller after averaging due to $x_{i}$ being an unbiased estimator.
For a training data $x$, a score-based model $s_{\theta}(x)$ can be trained to fit $\nabla_{x}\log p(x)$ for image generation. The distribution $x_{0}\sim p_{0}\approx\delta_{x}$ is a delta distribution supported on the trained image $x$.
Then for a given image $x$, we generate $n$ noisy images $\{y_{i}\}_{i\leq n}$, $y_{i}=\beta_{T}x+(1-\beta_{T})\epsilon_{i}$, and then denoise to get $n$ reconstructed image $\{\hat{x}_{i}\}_{i\leq n}$. If the score function $s_{\theta}(x)$ is learned properly, we will generate the image satisfying the following distribution:
$\hat{x}_{i}\sim\mathcal{N}(x,\sigma_{T}I),$
where $\sigma_{T}$ is a constant determined by the injected noise level $\beta_{T}$.
The key guarantee is that when the score function is learned for a data point $x$, then the reconstructed image $\hat{x}_{i}$ is an unbiased estimator of $x$. Hence, averaging over multiple independent samples $\hat{x}_{i}$ would greatly reduce the estimation error (see Theorem 1). On the other hand, for a non-member image $x^{\prime}$, the unbiasedness of the denoised image is not guaranteed. We illustrate the intuition in Figure 2.
With the above intuition, we introduce the details of our algorithm. We independently apply the black-box variation API $n$ times with our target image $x$ as input, average the output images, and then compare the average result $\hat{x}$ with the original image. We will discuss the impact of the averaging number $n$ in Section 6.4. We then evaluate the difference between the images using an indicator function:
$f(x)=\mathbf{1}[D(x,\hat{x})<\tau]\,.$
Our algorithm classifies a sample as being in the training set if $D(x,\hat{x})$ is smaller than a threshold $\tau$, where $D(x,\hat{x})$ represents the difference between the two images. It can be calculated using traditional functions, such as the SSIM metric [44]. Alternatively, we can train a neural network as a proxy. In Section 6, we will introduce the details of $D(x,\hat{x})$ used in our experiment.
Our algorithm is outlined in Algorithm 1, and we name it ReDiffuse. The key ideas of our algorithm are illustrated in Figure 1, and we also provide some theoretical analysis in Theorem 1 to support it.
Algorithm 1 Black-Box MIA through ReDiffuse
  Input: Target image $x$, diffusion step $t$, average number $n$, threshold $\tau$, the variation API of the target model $V_{\theta}$, distance function $D$.
  for $k=1,\ldots,n$ do
     Use the variation API $V_{\theta}$ to generate the variation image $\hat{x}_{k}=V_{\theta}(x,t)$ according to Equation (2).
  end for
  Average the reconstructed images from each iteration $\hat{x}=\frac{1}{n}(\hat{x}_{1}+\hat{x}_{2}+\ldots+\hat{x}_{n})$.
  return  "YES" if the distance between the two images $D(x,\hat{x})$ is less than $\tau$, otherwise "NO".
4.3 Theoretical Analysis
[…]
Theorem 1.
Suppose the DDIM model can learn a parameter $\theta$ such that, for any $x\sim D_{\text{training}}$, the prediction error $\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|$ is a centered random variable $X$ with finite cumulant-generating function [7]. We also assume the sampling interval $k$ is equal to the variation API diffusion step $t$. Let $\hat{x}$ be the average of $n$ output images of $x$ using the variation API. Then we have
$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq\exp(-n\Psi^{*}_{X}(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}))\,,$
where $\beta>0$ and $\Psi^{*}_{X}$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X}$.
The theorem suggests that averaging the randomly variational data from a training set will result in a smaller relative error with high probability when we use a large $n$. We defer the proof of this theorem to Appendix C.
5 MIA on Latent Diffusion Models
In this section, we discuss how we can generalize our algorithm to latent diffusion models. We will focus on stable diffusion [32], as it is one of the most popular choices.
5.1 The Variation API and Modification to ReDiffuse
We note that the variation API for stable diffusion is different from DDIM, as it includes the encoder-decoder process. Again we denote $\hat{x}=V_{\theta}(x,t)$, and the details are as follows:
$z=\text{Encoder}(x)\,,\quad z_{t}=\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\quad\hat{z}=\Phi_{\theta}(z_{t},0)\,,\quad\hat{x}=\text{Decoder}(\hat{z})\,.$ (4)
This definition aligns with the image generation process of the stable diffusion model.
Then we introduce the modification of our algorithm in this setup. For DDPM or DDIM, we evaluate the difference between the reconstructed image and the original image. However, for images processed by the stable diffusion model, which are typically larger in size, directly comparing the restored image with the original may lead to drastic differences. To enhance our algorithm, we modify the decision function, by independently adding random noise to the original image twice and then comparing the differences between the two restored images $\hat{x}_{1}$ and $\hat{x}_{2}$:
$f(x)=\mathbf{1}[D(\hat{x}_{1},\hat{x}_{2})<\tau]\,.$
This modification has shown promising experimental results, as detailed in Section 6.3. We refer to this improved version as “ReDiffuse+”.
5.2 LAION by DALL-E: A Style-shift Evaluation Dataset
Figure 3: Comparison of data used in our new experiment versus the previous work. It shows that our dataset, LAION by DALL-E, achieves minimal content differences from the original LAION-5B images. In contrast, the COCO dataset images used in previous works show significant content differences from the LAION-5B images, which could potentially impact the evaluation results.
The dataset used for evaluation of stable diffusion in prior studies [5, 18] possesses certain limitations. In the context of attacking stable diffusion, LAION-5B served as the in-training-set data, while the COCO dataset was designated as the out-of-training-set data. However, the inherent content disparities between these two datasets are considerable, shown in Figure 3. Therefore, the MIA algorithm’s capability to differentiate between them may not solely stem from whether the data was employed during training, but also from their content disparities.
Complementary to previous work, we propose a new evaluation dataset aimed at making the out-of-training-set data contain the same content as the training data but with different styles. Concretely, we use the textual descriptions from the LAION-5B dataset (which was used for training) to generate new images with the DALL-E model. We refer to this new dataset as “LAION by DALL-E” (shown in Figure 3). These generated images will serve as images that the stable diffusion model has not seen during training. By including this dataset in the evaluation, we achieve a more comprehensive evaluation of MIA methods for stable diffusion.
6 Experiments
In this section, we evaluate the performance of our methods across various datasets and settings. We follow the same experiment’s setup in previous papers [5, 18]. The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A.
6.1 Evaluation Metrics
We follow the metrics used in previous papers [5, 18], including Area Under Receiver Operating Characteristic (AUC), Attack Success Rate (ASR), the True Positive Rate (TP) when the False Positive Rate is 1%. We also plot the ROC curves.
6.2 MIA with DDIM Models
We train a DDIM model on the CIFAR-10/100 [20], STL10-Unlabeled [2] and Tiny-Imagenet [3] dataset, using the image generation step $T=1000$ and sampling interval $k=100$ (explained in Section 6.4). For all the datasets, we randomly select 50% of the training samples to train the model and denote them as members. The remaining 50% are utilized as nonmembers. We take [28], [5], [18] as our baseline method. We fix the diffusion step at $t=200$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$. We will discuss the impact of the diffusion step and the average number in Section 6.4.
Table 1: Comparison of different methods on four datasets for the DDIM model. We use AUC, ASR and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
Method CIFAR10 CIFAR100 STL10 TINY-IN
Algorithm U-Net AUC ASR TP AUC ASR TP AUC ASR TP AUC ASR TP
Loss [28] $\Box$ 0.88 0.82 14.2 0.92 0.84 20.9 0.89 0.82 15.6 0.94 0.87 29.4
SecMI [5] $\Box$ 0.95 0.90 40.7 0.96 0.90 44.9 0.94 0.88 26.9 0.97 0.94 48.4
PIA [18] $\Box$ 0.95 0.89 48.7 0.96 0.90 47.0 0.94 0.87 29.8 0.98 0.94 73.0
PIAN [18] $\Box$ 0.95 0.89 50.4 0.91 0.85 39.2 0.92 0.86 28.5 0.95 0.93 78.4
ReDiffuse $\blacksquare$ 0.96 0.91 40.7 0.98 0.93 48.2 0.96 0.90 31.9 0.99 0.96 78.8
$\Box$: Require the access of U-Net.  $\blacksquare$: Do not require the access of U-Net.
For the difference function $D(x,\hat{x})$, following the setup in [5], we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image. Using the ResNet-18 network [9] and denoting it as $f_{R}$, we perform binary classification on these difference vectors. We use 20% of the data as the training set and obtained the label of each difference vector being classified as a member or nonmember. The difference function is obtained by the negative value of the probability predicting as member by the neural network: $D(x,\hat{x})=-f_{R}(v)$.
The result is shown in Table 1. Our method achieve high performance, surpassing several baseline algorithms in most setups, and does not require access to the internal structure of the model. This demonstrates that our algorithm is highly effective and robust.
6.3 MIA with the Stable Diffusion Model
We conduct experiments on the original Stable Diffusion model, i.e., stable-diffusion-v1-5 provided by Huggingface, without further fine-tuning or modifications. First, we follow the experiment setup of [5, 18], use the LAION-5B dataset [36] as member and COCO2017-val [23] as non-member. We randomly select 2500 images in each dataset. We test two scenarios: Knowing the ground truth text, which we denote as Laion5 vs COCO; Generating text through BLIP [21], which we denote as Laion5 BLIP vs COCO.
Next, given the inherent content differences between the COCO dataset and the LAION dataset, as discussed in Section 4, we also include a new evaluation dataset. We randomly generated 500 images using DALL-E 3, using prompts derived from the captions associated with each image in LAION-5. We refer to this new dataset as “LAION by DALL-E”. This dataset serves as the non-member set, while the corresponding 500 images from LAION-5 serve as the member set. This evaluation focuses on style shift rather than content shift between the datasets, enabling a more comprehensive evaluation of Stable Diffusion.
For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric [44] to measure the differences between two images. Additionally, because the image sizes in these datasets are relatively large, the differences between the original and reconstructed images are generally significant. Therefore, besides the basic algorithm, we also propose a simple variant of our algorithm (named ReDiffuse+) based on methods discussed in Section 4, involving altering the image by API twice independently and directly measuring the differences between the two variational images, rather than comparing the variational images to the original images.
Table 2: Comparison of different methods for Stable Diffusion using the same set of metrics as Table 1. Again, previous methods require access to U-Net, whereas our methods do not.
Method Laion5 vs COCO Laion5 vs COCO with BLIP Laion5 vs Laion5 by DALL-E
AUC ASR TP AUC ASR TP AUC ASR TP
Loss [28] 0.62 0.61 13.2 0.62 0.62 13.3 0.52 0.55 7.6
SecMI [5] 0.70 0.65 19.2 0.71 0.66 19.8 0.66 0.64 15.2
PIA [18] 0.70 0.66 19.7 0.73 0.68 20.2 0.65 0.62 14.9
PIAN [18] 0.56 0.53 4.8 0.55 0.51 4.4 0.53 0.52 3.4
ReDiffuse 0.79 0.72 14.4 0.80 0.74 14.8 0.74 0.70 15.3
ReDiffuse+ 0.81 0.75 20.6 0.82 0.75 22.1 0.73 0.71 13.6
The results, presented in Table 2, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%. Notably, our methods do not require access to U-Net. In addition, previous methods perform poorly on the new evaluation dataset, failing to achieve effective membership inference attack detection. In contrast, our methods still achieve robust performance, indicating their ability to distinguish images in the training set even when differences in image content are excluded.
We also plot ROC curves for the Stable Diffusion model trained on Laion5 setups and the DDIM trained on CIFAR-10 in Figure 4. These curves further demonstrate the effectiveness of our algorithm.
Figure 4: The ROC curves of various methods trained on DDIM and Stable Diffusion. Left: Trained DDIM model on CIFAR-10. Right: Trained Stable Diffusion model on the LAION-5B dataset with the ground truth text from LAION-5B caption. Non-members are sampled from COCO. The curves show that our algorithm outperforms the baseline algorithms.
6.4 Ablation Studies
In this section, we alter some experimental parameters to test the robustness of our algorithm.
The Impact of Average Numbers
We use different average numbers $n$ and test the influence on the results. The results are shown in Figure 5. We observe that in the DDIM setup, due to the smaller image size in the dataset (32x32), there is a certain variance in the images obtained from each reconstruction. Averaging the images from multiple independent samples as the output $\hat{x}$ can further improve accuracy. In the stable diffusion setup, since the image size in the dataset is larger (512x512), the reconstructed images are more stable and not influenced by perturbations at specific coordinates. Therefore, averaging multiple images is not necessary. The figures of ASR are also plotted in Appendix B.
Figure 5: The impact of average numbers. Left: DDIM model on CIFAR-10. Averaging can further improve the performance of our algorithm. Right: Stable Diffusion model on LAION-5. In the latter case, averaging does not improve performance.
The Impact of Diffusion Steps
We adjust the diffusion step $t$ to examine its impact on the results. The experiments are conducted using the DDIM model trained on CIFAR-10 with diffusion steps for inference, ranging from 10 to 500, and the outcomes are presented in Figure 6. Our findings indicate that as long as a moderate step (50-300) is chosen, the attack performance remains excellent, demonstrating that our algorithm is not sensitive to the choice of $t$. This further underscores the robustness of our algorithm. We also plot the change of diffusion step for other datasets in the Appendix B.
Figure 6: The impact of diffusion steps. We train a DDIM model on the CIFAR-10 dataset and use different diffusion step for inference. We find that high accuracy can be achieved as long as a moderate step number is chosen. This opens up possibilities for practical applications in real-world scenarios.
The Impact of Sampling Intervals
In DDIM and the internal structure of Stable Diffusion, the model use a set of steps denoted by $\tau_{1},\tau_{2},\ldots,\tau_{T}$. It samples at each of these steps to create the image. The spacing between these steps is referred to as the sampling interval.
We change the sampling interval $k$ and check the influence on the results. As shown in Figure 7, we adjust this parameter for attacks on both DDIM and stable diffusion. We find that our method achieve high AUC values across different sampling intervals, demonstrating that our detection capabilities are not significantly limited by this parameter. Additionally, in Appendix B, we plot the effect of different sampling interval on ASR and find that the impact is minimal.
Figure 7: The impact of sampling intervals. Left: DDIM model on CIFAR-10. Right: Stable Diffusion model on LAION-5. We find that adjusting the sampling interval has a relatively small influence in the first case and does not affect our method in the latter case. This makes our algorithm applicable to more setups.
7 Conclusion, Limitations and Future Directions
In this work, we introduce a novel black-box membership inference attack method specifically designed for diffusion models. Our approach only requires access to the variation API of the model, bypassing the need for internal network components such as the U-net. This represents a significant advancement for commercial diffusion models, which typically restrict internal access. We demonstrate the effectiveness of our approach across various datasets, showing that it achieves high accuracy in identifying whether an image was included in the training dataset. Our method offers a practical solution for protecting the copyright of artworks.
However, our method has some limitations, particularly the need for a moderate diffusion step $t$ of the variation API. Our algorithm does not achieve high accuracy if the diffusion step is very high. Additionally, it would be desirable if we could expand the moderate size of our newly proposed evaluation dataset.
Future work could aim to derive more robust algorithms that can be used for a wider range of diffusion step. Improving the interpretability of our method and extending it to other generative models are also valuable directions for further research.
Appendix A Datasets, Models and Hyperparameters
We run experiments on DDIM and Stable Diffusion models, using an NVIDIA RTX 6000 graphics card for all experiments.
For DDIM, we follow the training hyperparameters of [5] to train a DDIM model on the CIFAR-10/100 [20], STL10-Unlabeled [2], and Tiny-Imagenet [3] datasets, using an image generation step of $T=1000$ and a sampling interval of $k=100$. The training iterations are set to 800,000. For all the datasets, we randomly select 50
We take [28], [5], and [18] as our baseline methods. We use their codes from the official repositories and apply the optimal hyperparameters from their papers. For our algorithm, we fix the diffusion step at $t=200$ and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
For the Stable Diffusion experiments, we use the original Stable Diffusion model, i.e., stable-diffusion-v1-5 provided by Huggingface, without further fine-tuning or modifications. We follow the experimental setup of [5, 18], using an image generation step of $T=1000$ and a sampling interval of $k=10$. We use the LAION-5B dataset [36] as the member set and COCO2017-val [23] as the non-member set. We randomly select 2500 images from each dataset. We test two scenarios: knowing the ground truth text, denoted as Laion5 vs. COCO; and generating text through BLIP [21], denoted as Laion5 BLIP vs. COCO. For the new dataset, we randomly generated 500 images using DALL-E 3, with prompts derived from the captions associated with each image in LAION-5. This dataset serves as the non-member set, while the corresponding 500 images from LAION-5 serve as the member set.
We use the hyperparameters from the papers [5, 18] to run the baseline methods. For our algorithms ReDiffuse  and ReDiffuse+, we fix the diffusion step at $t=10$ to call the variation API and directly use the SSIM metric [44] to measure the differences between two images. Other hyperparameters remain the same as in the baseline methods.
Appendix B More Experiment Results
In this section, we show other experiment results which is not in our main paper. We conduct more ablation studies.
The Impact of Average Numbers
We use different average numbers $n$ and test the influence on the results. Besides the figures of AUC in the main paper, the figures of ASR are also plotted in Figure 8. We observe that in the DDIM setup, averaging the images from multiple independent samples as the output further improves accuracy. In the Stable Diffusion setup, averaging multiple images is not necessary as it does not improve ASR performance.
Figure 8: The impact of average numbers. Left: DDIM model on CIFAR-10. Averaging can further improve the performance of our algorithm. Right: Stable Diffusion model on LAION-5. Averaging does not improve performance.
The Impact of Diffusion Steps
We adjust the diffusion step $t$ to examine its impact on the results. We plot the graphs for CIFAR100 (Figure 9), STL10 (Figure 10), and Laion5 vs. COCO (Figure 11). From the results, we see that our algorithm can achieve high performance over a wide range of diffusion steps, especially for Stable Diffusion. This opens up possibilities for practical applications in real-world scenarios.
Figure 9: The impact of diffusion steps. We train a DDIM model on the CIFAR-100 dataset and use different diffusion step for inference.
Figure 10: The impact of diffusion steps. We train a DDIM model on the STL-10 dataset and use different diffusion step for inference.
Figure 11: The impact of diffusion steps. We use the Stable Diffusion model with the Laion5 vs. COCO dataset for evaluation and test different diffusion steps for inference. Specifically, using relatively small diffusion steps results in better performance.
Impact of $L_{p}$ norm
For the DDIM model, we train a neural network using the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector v for each image. This corresponds to the $L_{1}$ norm. Then we change the $L_{p}$ norm of our neural network classifier to see if it has any influence on the results. The results are in Figure 12:
Figure 12: The impact of $L_{p}$ norms. Our algorithm remains effective for $p=1,2,3,4$.
From the results, we know that our algorithm is effective for $p=1,2,3,4$. We should not choose very large $p$ to train the neural network.
Impact of Sampling Intervals
We change the sampling intervals to see if there is any influence on the results. In addition to the AUC figures in the main paper, the ASR figures are also plotted in Figure 13. From the results, we observe that the different sampling intervals have minimal impact on ASR.
Figure 13: The impact of sampling intervals. Left: DDIM model on CIFAR-10. Right: Stable Diffusion model on LAION-5. We find that adjusting the sampling interval does not significantly affect the ASR.
Appendix C Proof
In this section, we implement the proof of the theorem. See 1
Proof.
We denote $x^{i}$ as the $i$-th output image of $x$ using the variation API. We denote the $i$-th Gaussian noise as $\epsilon^{i}$ and the forward process incur $x^{i}_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i}$.
As the sampling interval is equal to the variation API diffusion step $t$, we have
$x^{i}-x$ $=\frac{x^{i}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t)}{\sqrt{\bar{\alpha}_{t}}}-x\,,$
$=\frac{\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t)}{\sqrt{\bar{\alpha}_{t}}}-x\,,$
$=\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}(\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t))\,.$
If we denote the random variable $X_{i}$ represents $\|\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t))\|$, then we consider $S_{n}=\sum_{i=1}^{n}X_{i}$. From the assumption we know that all the $X_{i}$ are centered random variable $X$ with finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$. Using the theorem of Chernoff-Cramer method for sums of IID RV [7], we get the following probability inequality for any $\beta>0$:
$\mathbb{P}(S_{n}\geq\beta)\leq\exp(-n\Psi^{*}_{X}(\frac{\beta}{n}))\,.$
Therefore, taking the definition of $\|\hat{x}-x\|=\|\frac{1}{n}\sum_{i=1}^{n}(x^{i}-x)\|\leq\frac{\sqrt{1-\bar{\alpha}_{t}}}{n\sqrt{\bar{\alpha}_{t}}}S_{n}$, we get the following bound of the reconstruction error:
$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq\exp(-n\Psi^{*}_{X}(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}))\,,$
where $\beta>0$ and $\Psi^{*}_{X}(y)=\sup_{s>0}(sy-\Psi_{X}(s))$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X}$. So averaging the randomly reconstructed data from a training set will result in a smaller reconstruction error with high probability of $\Theta(1-\exp(-n))$ when we use a large $n$.

### v2

@@
 Abstract
-Identifying whether an artwork was used to train a diffusion model is an important research topic, given the rising popularity of AI-generated art and the associated copyright concerns.
+Given the rising popularity of AI-generated art and the associated copyright concerns, identifying whether an artwork was used to train a diffusion model is an important research topic.
 The work approaches this problem from the membership inference attack (MIA) perspective.
-We first identify the limitations of applying existing MIA methods for copyright protection: the required access of internal U-nets and the choice of non-member datasets for evaluation.
-To address the above problems, we introduce a novel black-box membership inference attack method that operates without needing access to the model’s internal U-net.
-We then construct a DALL-E generated dataset for a more comprehensive evaluation.
-We validate our method across various setups, and our experimental results outperform previous works.
+We first identify the limitation of applying existing MIA methods for proprietary diffusion models: the required access of internal U-nets.
+To address the above problem, we introduce a novel membership inference attack method that uses only the image-to-image variation API and operates without access to the model’s internal U-net.
+Our method is based on the intuition that the model can more easily obtain an unbiased noise prediction estimate for images from the training set.
+By applying the API multiple times to the target image, averaging the outputs, and comparing the result to the original image, our approach can classify whether a sample was part of the training set.
+We validate our method using DDIM and Stable Diffusion setups and further extend both our approach and existing algorithms to the Diffusion Transformer architecture.
+Our experimental results consistently outperform previous methods.
 1 Introduction
@@
 These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation [13, 43], text-to-image generation [35, 52, 30] and image-to-image generation [36].
-This surge has given rise to powerful AI art models such as DALL-E 2 [31], Stable Diffusion [32] and Imagen [34].
-The field of AI-generated art holds a promising future and is expected to have widespread impact.
+This surge has given rise to powerful AI art models such as DALL-E 2 [34], Stable Diffusion [35], and Imagen [37].
+AI-generated art holds a promising future and is expected to have a widespread impact.
 Effective training of diffusion models requires high-quality data.
 It is thus crucial to design an algorithm that can identify whether a specific artwork has been used during the training of a model, thereby providing protection for these artworks and detecting misuse of data.
+This is especially important due to the rapid growth of generative models, which has raised concerns over intellectual property (IP) rights, data privacy, and the ethical implications of training on copyrighted or proprietary content without consent.
+As these models are increasingly deployed across industries, detecting whether a specific piece of content was used in training can help prevent unauthorized use of artistic works, protecting creators’ copyrights and ownership rights.
 This is a classic problem in the field of machine learning, first introduced by [40] and named “membership inference attack”.
@@
 [6]and [19] extended this work by relaxing the assumptions about model access requirements.
-Despite great progress, previous methods are not yet ready for copyright protection in diffusion models for two reasons.
-First, most existing approaches heavily rely on checking whether the U-net of the model predicts noise accurately, which is not practical since most commercial diffusion models available today offer only API access, while the U-net remains hidden.
-We refer to MIA with access only to generated images as the Black-Box MIA.
-Second, existing evaluations for stable diffusion models primarily focus on content shifts in the data, while the shifts in art styles are not covered (illustrated in Figure 3).
-To address the above issues, we propose a black-box membership inference attack method that does not require access to the U-net of the model.
-We posit that if we alter an image using the target diffusion model’s variation API, the sampling process will be captured by a region of “attraction” if the model has seen this image during training (illustrated in Figure 2).
-By quantifying the disparity between the resulting reconstructed image with the original, we can effectively determine whether the image was part of the training dataset.
-We also construct a new evaluation dataset named LAION by DALL-E reflecting style shifts in data rather than content shifts.
-In addition to evaluating our methods on existing datasets across multiple models, we test MIA on the stable diffusion model to classify LAION by DALL-E from the ground truth training data LAION-5B.
-The main contributions of this work are listed as follows:
+Despite great progress, previous methods are not yet ready for MIA in proprietary diffusion models.
+Most existing approaches heavily rely on checking whether the U-net of the model predicts noise accurately, which is not practical since most commercial diffusion models available today offer only API access, while the U-net remains hidden.
+To address the above issue, we propose a membership inference attack method that relies on the variation API and does not require access to the denoise model (e.g., U-net).
+We observe that if we alter an image using the target diffusion model’s variation API, the sampling process will be captured by “a region of attraction” if the model has seen this image during training (illustrated in Figure 2).
+Based on the above observation, we propose the ReDiffuse algorithm for MIA with image-to-image variation API, and can detect member images without accessing the denoise model.
+Our main contributions are listed as follows:
   1. 1.
-    We propose a black-box membership inference attack method that does not require access to the model’s internal structure.
-Instead, we leverage the unique properties of diffusion models.
+    We propose a membership inference attack method that does not require access to the model’s internal structure.
 Our method only involves using the model’s variation API to alter an image and compare it with the original one.
@@
   2. 2.
-    We evaluate our method using DDIM and Stable Diffusion models on classical datasets, including CIFAR10, STL10-Unlabeled [2], LAION-5B [36], etc.
-Our method is shown to be effective and robust across datasets, and it outperforms the previous methods.
+    We evaluate our method using DDIM [43] and Stable Diffusion [35] models on classical datasets, including CIFAR10/100 [21], STL10-Unlabeled [2], LAION-5B [39], etc.
+Our method outperforms the previous methods.
   3. 3.
-    We propose a new MIA evaluation dataset which is complementary to previous work.
-We use the image titles from LAION-5B as prompts for DALL-E’s API [31] to generate images of the same contents but different styles.
-Under this setup, previous membership inference attack algorithms are no longer effective, while our method remains robust.
+    We extend both existing algorithms and our own algorithm to the Diffusion Transformer [32] architecture, implementing the membership inference attack within this model framework for the first time.
+Experimental results demonstrate that our algorithm is consistently effective.
 2 Related Works
@@
 Data Safety and Membership Inference Attack
-In the era of big data, preserving data privacy is paramount, especially in training diffusion models using sensitive datasets like artists’ artworks, which are protected by copyright laws.
+In the era of big data, preserving data privacy is paramount.
+The training of diffusion models may involve sensitive datasets like artists’ artworks, which are protected by copyright laws.
 Membership inference attacks, initially introduced by [40], serve as an effective means to detect potential misuse of data without proper authorization.
@@
 In the domain of membership inference attacks against diffusion models, [50, 15] use a white-box approach, which assumes access to the entire diffusion model and utilizes loss and likelihood to determine whether a sample is in the training set.
-More recently, [5, 18, 43] have relaxed these requirements, eliminating the need for the entire model.
+[6, 19, 46] have relaxed these requirements, eliminating the need for the entire model.
 They leverage the insight that samples within the training set yield more accurate noise predictions, thereby achieving high accuracy in membership inference attack tasks.
 However, they also require the outputs of the U-net, as it is necessary to obtain the noise predictions of intermediate steps.
+Recently, [31] proposed a black-box membership inference attack method against diffusion models.
+Their method identifies whether a specific image is in a finetuning dataset of 100 images by calculating the difference between the generated image and the target image, using the corresponding prompt as input.
+Their approach leverages the model’s tendency to memorize finetuning images and generate similar outputs.
+In contrast, we focus on detecting whether an image is in the pretraining dataset, where a pre-trained model often produces diverse outputs for the same prompt, making detection more challenging.
 3 Preliminary
-In this section, we begin by discussing several popular diffusion models.
+In this section, we begin by introducing the notations used for several popular diffusion models.
 We first introduce the Denoising Diffusion Probabilistic Model (DDPM) [13].
-Then, we extend to the Denoising Diffusion Implicit Model (DDIM) [40], which is a variant of DDPM used to accelerate image generation.
-Lastly, we discuss the Stable Diffusion [32], a model that condenses images into a latent space and generates images grounded in text descriptions.
+Then, we extend to the Denoising Diffusion Implicit Model (DDIM) [43] and Stable Diffusion [35], which are variants of DDPM used to accelerate image generation or generate images grounded in text descriptions.
+Lastly, we discuss Diffusion Transformer [32], a model that replaces the U-net architecture with a transformer and achieves higher-quality image generation.
 Denoising Diffusion Probabilistic Model (DDPM)
@@
 The forward process (denoted as $q$) iteratively incorporates Gaussian noise into an image, while the reverse process (denoted as $p_{\theta}$) gradually reconstructs the image from noise.
-$q(x_{t}\mid x_{t-1})$ $=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})\,,$
-$p_{\theta}(x_{t-1}\mid x_{t})$ $=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))\,,$
+$q(x_{t}\mid x_{t-1})$ $=\mathcal{N}\left(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}\right)\,,$
+$p_{\theta}(x_{t-1}\mid x_{t})$ $=\mathcal{N}\left(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)\right)\,,$
 where $\mu_{\theta}(\cdot)$ and $\Sigma_{\theta}(\cdot)$ are the mean and covariance of the denoised image parameterized by the model parameters $\theta$, and $\beta_{t}$ is a noise schedule that controls the amount of noise added at each step.
@@
 This approach significantly reduces computational costs and allows for high-quality image synthesis from textual descriptions.
+Diffusion Transformer
+Diffusion Transformer leverages the Vision Transformer [5] structure to replace the U-net architecture traditionally used in diffusion models for noise prediction.
+Its training and sampling methods remain consistent with DDIM, with the only difference being the replacement of noise prediction network $\epsilon_{\theta}$ with $\epsilon_{\tilde{\theta}}$, where $\tilde{\theta}$ represents a Vision Transformer-based architecture.
+This approach further enhances the generation quality and ensured that the model possesses good scalability properties.
 4 Algorithm Design
 In this section, we introduce our algorithm.
-We begin by discussing the definition of black-box API and the limitations of previous membership inference attack methods.
-In our formulations we assume DDIM as our target model.
-The formulations for DDPM is highly similar and we omit it for brevity.
-We will discuss the generalization to the latent diffusion model in Section 5.
-4.1 Towards Black-Box MIA for Diffusion Models
-Most previous works on membership inference attacks against diffusion models aim to prevent data leakage, and hence can rely on thresholding the model’s training loss.
+We begin by discussing the definition of variation API and the limitations of previous membership inference attack methods.
+In our formulations, we assume DDIM as our target model.
+The formulations for DDPM are highly similar and we omit it for brevity.
+We will discuss the generalization to the latent diffusion model in Section 4.3.
+4.1 The variation API for Diffusion Models
+Most previous works on membership inference attacks against diffusion models aim to prevent data leakage and hence rely on thresholding the model’s training loss.
 For instance, [15] involves a direct comparison of image losses, while [6, 19] evaluates the accuracy of the model’s noise prediction at initial or intermediate steps.
-However, the required access to the model’s internal U-net structure prevent applications to copyright protection because most servers typically provide only black-box API access.
-We denote a model API as black-box if it only requires the input image and returns the output image.
-Specifically, we do not have access to the internal structure of the model.
-Below we formulate the definition of the black-box variation API used in our algorithm.
-Definition 1 (The black-box variation API).
+However, the required access to the model’s internal U-net structure prevents applications from copyright protection because most servers typically provide only black-box API access.
+In contrast, our method represents a step towards black-box MIA, as we do not directly access the model’s internal structure.
+Instead, we rely solely on the variation API, which takes an input image and returns the corresponding output image.
+Below, we formalize the definition of the variation API used in our algorithm.
+Definition 1 (The variation API).
 We define the variation API $V_{\theta}(x,t)$ of a model as follows.
@@
 We will discuss the influence of different diffusion steps in Section 5.5, showing that the attack performances are relatively stable and not sensitive to the selection of $t$.
-We also note that the for the target model, we can substitute $\phi_{\theta}(x_{t},0)$ with other sampling methods, such as the Euler-Maruyama Method [27] or Variational Diffusion Models [16].
+We also note that for the target model, we can substitute $\phi_{\theta}(x_{t},0)$ with other sampling methods, such as the Euler-Maruyama Method [28] or Variational Diffusion Models [17].
 Figure 1: The overview of ReDiffuse.
-We independently input the image to the black-box variation API $n$ times with diffusion step $t$.
+We independently input the image to the variation API $n$ times with diffusion step $t$.
 We take the average of the output images and compare them with the original ones.
@@
 4.2 Algorithm
-In this section, we present our algorithm for black-box MIA.
-Our key intuition comes from the reverse SDE dynamics in continuous diffusion models [41, 42].
-Specifically, one simplified form of the reverse SDE (i.e., the denoise step) has the formulation below:
-$\dot{X_{t}}=(X_{t}/2-\nabla_{x}\log p(X_{t}))+dW_{t}\,,\quad t\in[0,T]\,.$ (3)
-where $W_{t}$ denotes a standard Brownian motion (with abuse of notation $t$) and $T$ is the generation time.
+[…]
+$L(\theta)=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right\|^{2}\right]\,.$
+Denote $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, we assume that the denoise model is expressive enough such that for the input $x_{0}\in\mathbb{R}^{d}$ and time step $t\in\mathcal{T}$, the Jacobian matrix $\nabla_{\theta}\epsilon_{\theta}(x_{t},t)$ is full rank $(\geq d)$.
+This suggest that the model can adjust the predicted noise $\epsilon_{\theta}(x_{t},t)$ locally in any direction.
+Then for a well trained model, we would have
+$\nabla_{\theta}L(\theta)=0$ $\implies\nabla_{\theta}\epsilon_{\theta}(x_{t},t)^{T}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right]=0\,,$
+$\implies\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right]=0\,.$
+Intuitively, this is because if the noise prediction from the neural network exhibited high bias, the network could adjust to fit the bias term, further reducing the training loss.
 Figure 2: The intuition of our algorithm design.
-We denote $x$ as the target image, $x_{i}$ as the $i$-th image generated by the variation API, and $\hat{x}$ as the average image of them.
+We denote $x$ as the target image, $\hat{x}_{i}$ as the $i$-th image generated by the variation API, and $\hat{x}$ as the average image of them.
 For member image $x$, the difference $v=x-\hat{x}$ will be smaller after averaging due to $x_{i}$ being an unbiased estimator.
-For a training data $x$, a score-based model $s_{\theta}(x)$ can be trained to fit $\nabla_{x}\log p(x)$ for image generation.
-The distribution $x_{0}\sim p_{0}\approx\delta_{x}$ is a delta distribution supported on the trained image $x$.
-Then for a given image $x$, we generate $n$ noisy images $\{y_{i}\}_{i\leq n}$, $y_{i}=\beta_{T}x+(1-\beta_{T})\epsilon_{i}$, and then denoise to get $n$ reconstructed image $\{\hat{x}_{i}\}_{i\leq n}$.
-If the score function $s_{\theta}(x)$ is learned properly, we will generate the image satisfying the following distribution:
-$\hat{x}_{i}\sim\mathcal{N}(x,\sigma_{T}I),$
-where $\sigma_{T}$ is a constant determined by the injected noise level $\beta_{T}$.
-The key guarantee is that when the score function is learned for a data point $x$, then the reconstructed image $\hat{x}_{i}$ is an unbiased estimator of $x$.
-Hence, averaging over multiple independent samples $\hat{x}_{i}$ would greatly reduce the estimation error (see Theorem 1).
-On the other hand, for a non-member image $x^{\prime}$, the unbiasedness of the denoised image is not guaranteed.
+Therefore, for images in the training set, we expect the network to provide an unbiased noise prediction.
+Since the noise prediction is typically inaccessible in practical applications, we use the reconstructed sample $\hat{x}$ as a proxy.
+Leveraging the unbiasedness of noise prediction, we demonstrate that averaging over multiple independent reconstructed samples $\hat{x}_{i}$ significantly reduces estimation error (see Theorem 1).
+On the other hand, for images that are not in the training set, the neural network may not provide an unbiased prediction at these points.
 We illustrate the intuition in Figure 2.
 With the above intuition, we introduce the details of our algorithm.
-We independently apply the black-box variation API $n$ times with our target image $x$ as input, average the output images, and then compare the average result $\hat{x}$ with the original image.
+We independently apply the variation API $n$ times with our target image $x$ as input, average the output images, and then compare the average result $\hat{x}$ with the original image.
 We will discuss the impact of the averaging number $n$ in Section 5.5.
 We then evaluate the difference between the images using an indicator function:
-$f(x)=\mathbf{1}[D(x,\hat{x})<\tau]\,.$
+$f(x)=\mathbf{1}\left[D(x,\hat{x})<\tau\right]\,.$
 Our algorithm classifies a sample as being in the training set if $D(x,\hat{x})$ is smaller than a threshold $\tau$, where $D(x,\hat{x})$ represents the difference between the two images.
@@
 The key ideas of our algorithm are illustrated in Figure 1, and we also provide some theoretical analysis in Theorem 1 to support it.
-Algorithm 1 Black-Box MIA through ReDiffuse
+Algorithm 1 MIA through ReDiffuse
   Input: Target image $x$, diffusion step $t$, average number $n$, threshold $\tau$, the variation API of the target model $V_{\theta}$, distance function $D$.
@@
   return  "YES" if the distance between the two images $D(x,\hat{x})$ is less than $\tau$, otherwise "NO".
-4.3 Theoretical Analysis
-[…]
+Analysis
+We give a short analysis to justify why averaging over $n$ samples in ReDiffuse can reduce the prediction error for training data.
+We have the following theorem showing that if we use the variation API to input a member $x\sim D_{\text{training}}$, then the error $\|\hat{x}-x\|$ from our method will be small with high probability.
 Theorem 1.
-Suppose the DDIM model can learn a parameter $\theta$ such that, for any $x\sim D_{\text{training}}$, the prediction error $\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|$ is a centered random variable $X$ with finite cumulant-generating function [7].
-We also assume the sampling interval $k$ is equal to the variation API diffusion step $t$.
+Suppose the DDIM model can learn a parameter $\theta$ such that, for any $x\sim D_{\text{training}}$ with dimension $d$, the prediction error $\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)$ is a random variable $X=(X_{1},X_{2},\ldots,X_{d})$ with zero expectation and finite cumulant-generating function for each coordinate [8].
+Suppose the sampling interval $k$ is equal to the variation API diffusion step $t$.
 Let $\hat{x}$ be the average of $n$ output images of $x$ using the variation API.
 Then we have
-$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq\exp(-n\Psi^{*}_{X}(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}))\,,$
-where $\beta>0$ and $\Psi^{*}_{X}$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X}$.
+$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq d\exp\left(-n\min_{i}\Psi^{*}_{X_{i}}\left(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{d(1-\bar{\alpha}_{t})}}\right)\right)\,,$
+where $\beta>0$ and $\Psi^{*}_{X_{i}}$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X_{i}}$.
 The theorem suggests that averaging the randomly variational data from a training set will result in a smaller relative error with high probability when we use a large $n$.
 We defer the proof of this theorem to Appendix C.
-5 MIA on Latent Diffusion Models
-In this section, we discuss how we can generalize our algorithm to latent diffusion models. We will focus on stable diffusion [32], as it is one of the most popular choices.
-5.1 The Variation API and Modification to ReDiffuse
-We note that the variation API for stable diffusion is different from DDIM, as it includes the encoder-decoder process. Again we denote $\hat{x}=V_{\theta}(x,t)$, and the details are as follows:
+We note that the unbiased assumption on predicted noise can be strong in practice.
+For experiments, we expect the mean $\|\mu\|$ of predicted noise for trained data is smaller than the mean $\|\mu^{\prime}\|$ from unseen data.
+As a result, the empirical best choice of $n$ would be determined by the gap $\|\mu^{\prime}\|-\|\mu\|$.
+4.3 MIA on Other Diffusion Models
+In this section, we discuss how we can generalize our algorithm to other diffusion models.
+We note that the variation API for stable diffusion is different from DDIM, as it includes the encoder-decoder process.
+Again we denote $\hat{x}=V_{\theta}(x,t)$, and the details are as follows:
 $z=\text{Encoder}(x)\,,\quad z_{t}=\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\quad\hat{z}=\Phi_{\theta}(z_{t},0)\,,\quad\hat{x}=\text{Decoder}(\hat{z})\,.$ (3)
 This definition aligns with the image generation process of the stable diffusion model.
-Then we introduce the modification of our algorithm in this setup.
-For DDPM or DDIM, we evaluate the difference between the reconstructed image and the original image.
-However, for images processed by the stable diffusion model, which are typically larger in size, directly comparing the restored image with the original may lead to drastic differences.
-To enhance our algorithm, we modify the decision function, by independently adding random noise to the original image twice and then comparing the differences between the two restored images $\hat{x}_{1}$ and $\hat{x}_{2}$:
-$f(x)=\mathbf{1}[D(\hat{x}_{1},\hat{x}_{2})<\tau]\,.$
-This modification has shown promising experimental results, as detailed in Section 6.3. We refer to this improved version as “ReDiffuse+”.
-5.2 LAION by DALL-E: A Style-shift Evaluation Dataset
-Figure 3: Comparison of data used in our new experiment versus the previous work.
-It shows that our dataset, LAION by DALL-E, achieves minimal content differences from the original LAION-5B images.
-In contrast, the COCO dataset images used in previous works show significant content differences from the LAION-5B images, which could potentially impact the evaluation results.
-The dataset used for evaluation of stable diffusion in prior studies [5, 18] possesses certain limitations.
-In the context of attacking stable diffusion, LAION-5B served as the in-training-set data, while the COCO dataset was designated as the out-of-training-set data.
-However, the inherent content disparities between these two datasets are considerable, shown in Figure 3.
-Therefore, the MIA algorithm’s capability to differentiate between them may not solely stem from whether the data was employed during training, but also from their content disparities.
-Complementary to previous work, we propose a new evaluation dataset aimed at making the out-of-training-set data contain the same content as the training data but with different styles.
-Concretely, we use the textual descriptions from the LAION-5B dataset (which was used for training) to generate new images with the DALL-E model.
-We refer to this new dataset as “LAION by DALL-E” (shown in Figure 3).
-These generated images will serve as images that the stable diffusion model has not seen during training.
-By including this dataset in the evaluation, we achieve a more comprehensive evaluation of MIA methods for stable diffusion.
+For the Diffusion Transformer, we define the variation API as $V_{\tilde{\theta}}(x,t)$, where $\tilde{\theta}$ corresponds to the Vision Transformer architecture instead of the U-net.
+We repeatedly call the variation API and calculate the difference between the original image and the reconstructed image, as done in DDIM.
 5 Experiments
 In this section, we evaluate the performance of our methods across various datasets and settings.
-We follow the same experiment’s setup in previous papers [5, 18].
+We follow the same experiment setup in previous papers [6, 19].
 The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A.
@@
 5.2 MIA with DDIM Models
-We train a DDIM model on the CIFAR-10/100 [20], STL10-Unlabeled [2] and Tiny-Imagenet [3] dataset, using the image generation step $T=1000$ and sampling interval $k=100$ (explained in Section 6.4).
+We follow the experimental setup of [6, 19].
+We train a DDIM model on the CIFAR-10/100 [21] and STL10-Unlabeled datasets [2], using the image generation step $T=1000$ and sampling interval $k=100$.
 For all the datasets, we randomly select 50% of the training samples to train the model and denote them as members.
 The remaining 50% are utilized as nonmembers.
-We take [28], [5], [18] as our baseline method.
+We use [29], [6], [19] as our baseline methods.
 We fix the diffusion step at $t=200$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
 We will discuss the impact of the diffusion step and the average number in Section 5.5.
-Table 1: Comparison of different methods on four datasets for the DDIM model. We use AUC, ASR and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
-Method CIFAR10 CIFAR100 STL10 TINY-IN
-Algorithm U-Net AUC ASR TP AUC ASR TP AUC ASR TP AUC ASR TP
-Loss [28] $\Box$ 0.88 0.82 14.2 0.92 0.84 20.9 0.89 0.82 15.6 0.94 0.87 29.4
-SecMI [5] $\Box$ 0.95 0.90 40.7 0.96 0.90 44.9 0.94 0.88 26.9 0.97 0.94 48.4
-PIA [18] $\Box$ 0.95 0.89 48.7 0.96 0.90 47.0 0.94 0.87 29.8 0.98 0.94 73.0
-PIAN [18] $\Box$ 0.95 0.89 50.4 0.91 0.85 39.2 0.92 0.86 28.5 0.95 0.93 78.4
-ReDiffuse $\blacksquare$ 0.96 0.91 40.7 0.98 0.93 48.2 0.96 0.90 31.9 0.99 0.96 78.8
+Table 1: Comparison of different methods on four datasets for the DDIM model. We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
+Method CIFAR10 CIFAR100 STL10
+Algorithm U-Net AUC ASR TP AUC ASR TP AUC ASR TP
+Loss [29] $\Box$ 0.88 0.82 14.2 0.92 0.84 20.9 0.89 0.82 15.6
+SecMI [6] $\Box$ 0.95 0.90 40.7 0.96 0.90 44.9 0.94 0.88 26.9
+PIA [19] $\Box$ 0.95 0.89 48.7 0.96 0.90 47.0 0.94 0.87 29.8
+PIAN [19] $\Box$ 0.95 0.89 50.4 0.91 0.85 39.2 0.92 0.86 28.5
+ReDiffuse $\blacksquare$ 0.96 0.91 40.7 0.98 0.93 48.2 0.96 0.90 31.9
 $\Box$: Require the access of U-Net.  $\blacksquare$: Do not require the access of U-Net.
@@
 We use 20% of the data as the training set and obtained the label of each difference vector being classified as a member or nonmember.
-The difference function is obtained by the negative value of the probability predicting as member by the neural network: $D(x,\hat{x})=-f_{R}(v)$.
+The difference function is obtained by the negated value of the probability outputed by the neural network predicting as member: $D(x,\hat{x})=-f_{R}(v)$.
 The result is shown in Table 1.
-Our method achieve high performance, surpassing several baseline algorithms in most setups, and does not require access to the internal structure of the model.
+Our method achieves high performance, surpassing several baseline algorithms in most setups, and does not require access to the internal structure of the model.
 This demonstrates that our algorithm is highly effective and robust.
+5.3 MIA with Diffusion Transformers
+We train a diffusion transformer model on the ImageNet [3] dataset following the setup of [32].
+We randomly select 100,000 images from the ImageNet training set to train the model with resolutions of either $128\times 128$ or $256\times 256$.
+For the membership inference attack setup, 1000 images are randomly chosen from our training set as the member set, and another 1000 images are randomly selected from the ImageNet validation set as the non-member set.
+We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
+We use [29], [6], [19] as our baseline methods.
+Since these work did not study the case of Diffusion Transformers, we integrate their algorithms into the DiT framework for evaluation.
+For the difference function $D(x,\hat{x})$, following the setup in [6, 19], we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
+Table 2: Comparison of different methods for Diffusion Transformer using the same set of metrics as Table 1.
+Previous methods require access to the Vision Transformer, whereas our methods do not.
+We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
+Method ImageNet 128 $\times$ 128 ImageNet 256 $\times$ 256
+Algorithm Transformer AUC ASR TP AUC ASR TP
+Loss [29] $\Box$ 0.83 0.76 10.7 0.78 0.70 7.3
+SecMI [6] $\Box$ 0.80 0.73 8.3 0.88 0.80 16.3
+PIA [19] $\Box$ 0.97 0.92 32.1 0.91 0.85 6.8
+PIAN [19] $\Box$ 0.66 0.64 6.2 0.67 0.66 12.8
+ReDiffuse $\blacksquare$ 0.98 0.95 44.1 0.97 0.94 47.3
+$\Box$: Require the access of Transformer.  $\blacksquare$: Do not require the access of Transformer.
+We also plot ROC curves for the DDIM train on CIFAR-10 and the Diffusion Transformer train on ImageNet $256\times 256$ in Figure 3. The curves further demonstrate the effectiveness of our method.
+Figure 3: The ROC curves of various setups.
+Left: DDIM model on CIFAR-10.
+Right: Diffusion Transformer on ImageNet $256\times 256$.
+The curves show that our algorithm outperforms the baseline algorithms.
 5.4 MIA with the Stable Diffusion Model
-We conduct experiments on the original Stable Diffusion model, i.e., stable-diffusion-v1-5 provided by Huggingface, without further fine-tuning or modifications.
-First, we follow the experiment setup of [5, 18], use the LAION-5B dataset [36] as member and COCO2017-val [23] as non-member.
+We conduct experiments on the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications.
+We follow the experiment setup of [6, 19], use the LAION-5B dataset [39] as member and COCO2017-val [24] as non-member.
 We randomly select 2500 images in each dataset.
-We test two scenarios: Knowing the ground truth text, which we denote as Laion5 vs COCO; Generating text through BLIP [21], which we denote as Laion5 BLIP vs COCO.
-Next, given the inherent content differences between the COCO dataset and the LAION dataset, as discussed in Section 4, we also include a new evaluation dataset.
-We randomly generated 500 images using DALL-E 3, using prompts derived from the captions associated with each image in LAION-5.
-We refer to this new dataset as “LAION by DALL-E”.
-This dataset serves as the non-member set, while the corresponding 500 images from LAION-5 serve as the member set.
-This evaluation focuses on style shift rather than content shift between the datasets, enabling a more comprehensive evaluation of Stable Diffusion.
+We test two scenarios: Knowing the ground truth text, which we denote as Laion5; Not knowing the ground truth text and generating text through BLIP [22], which we denote as Laion5 with BLIP.
 For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric [47] to measure the differences between two images.
-Additionally, because the image sizes in these datasets are relatively large, the differences between the original and reconstructed images are generally significant.
-Therefore, besides the basic algorithm, we also propose a simple variant of our algorithm (named ReDiffuse+) based on methods discussed in Section 4, involving altering the image by API twice independently and directly measuring the differences between the two variational images, rather than comparing the variational images to the original images.
-Table 2: Comparison of different methods for Stable Diffusion using the same set of metrics as Table 1. Again, previous methods require access to U-Net, whereas our methods do not.
-Method Laion5 vs COCO Laion5 vs COCO with BLIP Laion5 vs Laion5 by DALL-E
-AUC ASR TP AUC ASR TP AUC ASR TP
-Loss [28] 0.62 0.61 13.2 0.62 0.62 13.3 0.52 0.55 7.6
-SecMI [5] 0.70 0.65 19.2 0.71 0.66 19.8 0.66 0.64 15.2
-PIA [18] 0.70 0.66 19.7 0.73 0.68 20.2 0.65 0.62 14.9
-PIAN [18] 0.56 0.53 4.8 0.55 0.51 4.4 0.53 0.52 3.4
-ReDiffuse 0.79 0.72 14.4 0.80 0.74 14.8 0.74 0.70 15.3
-ReDiffuse+ 0.81 0.75 20.6 0.82 0.75 22.1 0.73 0.71 13.6
 The results, presented in Table 3, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%.
 Notably, our methods do not require access to U-Net.
-In addition, previous methods perform poorly on the new evaluation dataset, failing to achieve effective membership inference attack detection.
-In contrast, our methods still achieve robust performance, indicating their ability to distinguish images in the training set even when differences in image content are excluded.
-We also plot ROC curves for the Stable Diffusion model trained on Laion5 setups and the DDIM trained on CIFAR-10 in Figure 4. These curves further demonstrate the effectiveness of our algorithm.
-Figure 4: The ROC curves of various methods trained on DDIM and Stable Diffusion.
-Left: Trained DDIM model on CIFAR-10.
-Right: Trained Stable Diffusion model on the LAION-5B dataset with the ground truth text from LAION-5B caption.
-Non-members are sampled from COCO.
-The curves show that our algorithm outperforms the baseline algorithms.
+Table 3: Comparison of different methods for Stable Diffusion using the same set of metrics as Table 1.
+Again, previous methods require access to U-Net, whereas our methods do not.
+We use AUC, ASR and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
+Method Laion5 Laion5 with BLIP
+Algorithm U-Net AUC ASR TP AUC ASR TP
+Loss [29] $\Box$ 0.62 0.61 13.2 0.62 0.62 13.3
+SecMI [6] $\Box$ 0.70 0.65 19.2 0.71 0.66 19.8
+PIA [19] $\Box$ 0.70 0.66 19.7 0.73 0.68 20.2
+PIAN [19] $\Box$ 0.56 0.53 4.8 0.55 0.51 4.4
+ReDiffuse $\blacksquare$ 0.81 0.75 20.6 0.82 0.75 21.7
+$\Box$: Require the access of U-Net.  $\blacksquare$: Do not require the access of U-Net.
 5.5 Ablation Studies
 In this section, we alter some experimental parameters to test the robustness of our algorithm.
+We primarily focus on the ablation study of DDIM and Diffusion Transformer, while the ablation study related to Stable Diffusion is provided in Appendix B.
 The Impact of Average Numbers
-We use different average numbers $n$ and test the influence on the results.
-The results are shown in Figure 5.
-We observe that in the DDIM setup, due to the smaller image size in the dataset (32x32), there is a certain variance in the images obtained from each reconstruction.
-Averaging the images from multiple independent samples as the output $\hat{x}$ can further improve accuracy.
-In the stable diffusion setup, since the image size in the dataset is larger (512x512), the reconstructed images are more stable and not influenced by perturbations at specific coordinates.
-Therefore, averaging multiple images is not necessary.
-The figures of ASR are also plotted in Appendix B.
+We test the effect of using different averaging numbers $n$ on the results, as shown in Figure 4 .
+It can be observed that averaging the images from multiple independent samples to generate the output $\hat{x}$ further improves accuracy.
+This observation validates the algorithm design intuition discussed in Section 4.2.
+Additional figures showing the ASR results are presented in Appendix B.
 Figure 4: The impact of average numbers.
 Left: DDIM model on CIFAR-10.
-Averaging can further improve the performance of our algorithm.
-Right: Stable Diffusion model on LAION-5.
-In the latter case, averaging does not improve performance.
+Right: Diffusion Transformer model on Imagenet.
+In both cases, averaging multiple independent samples proves to be effective in further improving the overall performance of our algorithm, which validates the intuition of our algorithm design.
 The Impact of Diffusion Steps
 We adjust the diffusion step $t$ to examine its impact on the results.
-The experiments are conducted using the DDIM model trained on CIFAR-10 with diffusion steps for inference, ranging from 10 to 500, and the outcomes are presented in Figure 6.
-Our findings indicate that as long as a moderate step (50-300) is chosen, the attack performance remains excellent, demonstrating that our algorithm is not sensitive to the choice of $t$.
+The experiments are conducted using the DDIM model on CIFAR-10 with diffusion steps for inference.
+The outcomes are presented in Figure 5.
+Our findings indicate that as long as a moderate step is chosen, the attack performance remains excellent, demonstrating that our algorithm is not sensitive to the choice of $t$.
 This further underscores the robustness of our algorithm.
-We also plot the change of diffusion step for other datasets in the Appendix B.
-Figure 6: The impact of diffusion steps.
-We train a DDIM model on the CIFAR-10 dataset and use different diffusion step for inference.
+We also plot the change of diffusion step for other diffusion models and datasets in the Appendix B.
+Figure 5: The impact of diffusion steps on DDIM.
+We train a DDIM model on the CIFAR-10 dataset and use different diffusion steps for inference.
 We find that high accuracy can be achieved as long as a moderate step number is chosen.
@@
 The Impact of Sampling Intervals
-In DDIM and the internal structure of Stable Diffusion, the model use a set of steps denoted by $\tau_{1},\tau_{2},\ldots,\tau_{T}$.
-It samples at each of these steps to create the image.
+In DDIM and Diffusion Transformer, the model uses a set of steps denoted by $\tau_{1},\tau_{2},\ldots,\tau_{T}$.
+It samples each of these steps to create the image.
 The spacing between these steps is referred to as the sampling interval.
 We change the sampling interval $k$ and check the influence on the results.
-As shown in Figure 7, we adjust this parameter for attacks on both DDIM and stable diffusion.
-We find that our method achieve high AUC values across different sampling intervals, demonstrating that our detection capabilities are not significantly limited by this parameter.
-Additionally, in Appendix B, we plot the effect of different sampling interval on ASR and find that the impact is minimal.
+As shown in Figure 6, we adjust this parameter for attacks on both DDIM and Diffusion Transformer.
+We find that our method achieves high AUC values across different sampling intervals, demonstrating that our detection capabilities are not significantly limited by this parameter.
+Additionally, in Appendix B, we plot the effect of different sampling intervals on ASR and find that the impact is minimal.
 Figure 6: The impact of sampling intervals.
@@
 6 Conclusion, Limitations and Future Directions
-In this work, we introduce a novel black-box membership inference attack method specifically designed for diffusion models.
+In this work, we introduce a novel membership inference attack method specifically designed for diffusion models.
 Our approach only requires access to the variation API of the model, bypassing the need for internal network components such as the U-net.
-This represents a significant advancement for commercial diffusion models, which typically restrict internal access.
+This represents an advancement for commercial diffusion models, which typically restrict internal access.
 We demonstrate the effectiveness of our approach across various datasets, showing that it achieves high accuracy in identifying whether an image was included in the training dataset.
-Our method offers a practical solution for protecting the copyright of artworks.
-However, our method has some limitations, particularly the need for a moderate diffusion step $t$ of the variation API.
-Our algorithm does not achieve high accuracy if the diffusion step is very high.
-Additionally, it would be desirable if we could expand the moderate size of our newly proposed evaluation dataset.
-Future work could aim to derive more robust algorithms that can be used for a wider range of diffusion step.
+Our algorithm can detect data misuse by the model, representing a step forward in protecting the copyright of artworks.
+However, our method has certain limitations, particularly the requirement for a moderate diffusion step $t$ in the variation API.
+The algorithm’s accuracy declines when the diffusion step is excessively high.
+As such, we propose our method as an initial step towards black-box MIA, with a more comprehensive solution left for subsequent exploration.
+Future work could focus on developing more robust algorithms capable of handling a broader range of diffusion steps.
 Improving the interpretability of our method and extending it to other generative models are also valuable directions for further research.
 Appendix A Datasets, Models and Hyperparameters
-We run experiments on DDIM and Stable Diffusion models, using an NVIDIA RTX 6000 graphics card for all experiments.
-For DDIM, we follow the training hyperparameters of [5] to train a DDIM model on the CIFAR-10/100 [20], STL10-Unlabeled [2], and Tiny-Imagenet [3] datasets, using an image generation step of $T=1000$ and a sampling interval of $k=100$.
+We use NVIDIA RTX 6000 graphics cards for all our experiments.
+For DDIM, we follow the training hyperparameters of [6] to train a DDIM model on the CIFAR-10/100 [21] and STL10-Unlabeled [2] datasets, using 1000 image generation steps ($T=1000$) and a sampling interval of $k=100$.
 The training iterations are set to 800,000.
-For all the datasets, we randomly select 50
-We take [28], [5], and [18] as our baseline methods.
-We use their codes from the official repositories and apply the optimal hyperparameters from their papers.
-For our algorithm, we fix the diffusion step at $t=200$ and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
-For the Stable Diffusion experiments, we use the original Stable Diffusion model, i.e., stable-diffusion-v1-5 provided by Huggingface, without further fine-tuning or modifications.
+For all the datasets, we randomly select 50% of the training samples to train the model and designate them as members.
+The remaining 50% of the training samples are used as nonmembers.
+We use [29, 6, 19] as our baseline methods.
+We use their official code repositories and apply the optimal hyperparameters from their papers.
+For our algorithm, we fix the diffusion step at $t=200$ and independently call the variation API 10 times to average the output images as $\hat{x}$.
+For the Diffusion Transformers, we train the model on the ImageNet [3] dataset, using 1000 image generation steps ($T=1000$), while for other training hyperparameters, we follow the setup of [32].
+We randomly select 100,000 images from the ImageNet training set to train the model at resolutions of either $128\times 128$ or $256\times 256$.
+For the $128\times 128$ image size, we use 160,000 training iterations.
+For the $256\times 256$ image size, we use 300,000 training iterations.
+These numbers of training iterations are chosen to ensure the generation of high-quality images.
+For the membership inference attack setup, 1000 images are randomly selected from our training set as the member set, and another 1000 images are randomly chosen from the ImageNet validation set as the non-member set.
+We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to average the output images as $\hat{x}$.
+We also use [29, 6, 19] as our baseline methods.
+Since these papers do not study Diffusion Transformers, we adapt their algorithms to the DiT framework for evaluation.
+We fix the DDIM step at $k=50$ and choose diffusion steps $t\in[50,100,150,200,250,300]$, recording their optimal solutions under these different hyperparameter settings.
+For the Stable Diffusion experiments, we use the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications.
 We follow the experimental setup of [6, 19], using an image generation step of $T=1000$ and a sampling interval of $k=10$.
@@
 We randomly select 2500 images from each dataset.
-We test two scenarios: knowing the ground truth text, denoted as Laion5 vs.
-COCO; and generating text through BLIP [21], denoted as Laion5 BLIP vs.
-COCO.
-For the new dataset, we randomly generated 500 images using DALL-E 3, with prompts derived from the captions associated with each image in LAION-5.
-This dataset serves as the non-member set, while the corresponding 500 images from LAION-5 serve as the member set.
+We test two scenarios: knowing the ground truth text, denoted as Laion; and generating text through BLIP [22], denoted as Laion with BLIP.
 We use the hyperparameters from the papers [6, 19] to run the baseline methods.
-For our algorithms ReDiffuse  and ReDiffuse+, we fix the diffusion step at $t=10$ to call the variation API and directly use the SSIM metric [44] to measure the differences between two images.
+For our algorithms ReDiffuse, we fix the diffusion step at $t=10$ to call the variation API and directly use the SSIM metric [47] to measure the differences between two images.
 Other hyperparameters remain the same as in the baseline methods.
@@
 We use different average numbers $n$ and test the influence on the results.
-Besides the figures of AUC in the main paper, the figures of ASR are also plotted in Figure 8.
-We observe that in the DDIM setup, averaging the images from multiple independent samples as the output further improves accuracy.
-In the Stable Diffusion setup, averaging multiple images is not necessary as it does not improve ASR performance.
-Figure 8: The impact of average numbers.
-Left: DDIM model on CIFAR-10.
-Averaging can further improve the performance of our algorithm.
-Right: Stable Diffusion model on LAION-5.
-Averaging does not improve performance.
+Besides the figures of AUC in the main paper, the figures of ASR of DDIM and Diffusion Transformer are also plotted in Figure 7.
+In addition, we plot the figures of AUC and ASR for Stable Diffusion in Figure 8.
+We observe that in the DDIM and Diffusion Transformer setup, averaging the images from multiple independent samples as the output further improves accuracy.
+In the stable diffusion setup, since the image size in the dataset is larger (512x512), the reconstructed images are more stable and not influenced by perturbations at specific coordinates.
+Therefore, averaging multiple images is not necessary.
+Figure 7: The impact of average numbers. Left: DDIM model on CIFAR-10. Right: Diffusion Transformer model on Imagenet. Averaging can further improve the performance of our algorithm.
+Figure 8: The impact of average numbers on Stable Diffusion. We plot the AUC and ASR metrics, and averaging does not improve performance.
 The Impact of Diffusion Steps
 We adjust the diffusion step $t$ to examine its impact on the results.
-We plot the graphs for CIFAR100 (Figure 9), STL10 (Figure 10), and Laion5 vs.
-COCO (Figure 11).
-From the results, we see that our algorithm can achieve high performance over a wide range of diffusion steps, especially for Stable Diffusion.
+We train the DDIM model on CIFAR100 (Figure 9), STL10 (Figure 10) dataset, Diffusion Transformer on the Imagenet $256\times 256$ (Figure 11)dataset and Stable Diffusion on Laion5 dataset (Figure 12).
+From the results, we see that our algorithm can achieve high performance over a wide range of diffusion steps.
 This opens up possibilities for practical applications in real-world scenarios.
@@
 Figure 11: The impact of diffusion steps.
-We use the Stable Diffusion model with the Laion5 vs.
-COCO dataset for evaluation and test different diffusion steps for inference.
+We train a Diffusion Transformer model on the Imagenet $256\times 256$ dataset and use different diffusion step for inference.
+The robust results imply that our algorithm is also not very sensitive to the choice of $t$ in this setup.
+Figure 12: The impact of diffusion steps.
+We use the Stable Diffusion model with the Laion5 dataset for evaluation and test different diffusion steps for inference.
 Specifically, using relatively small diffusion steps results in better performance.
-Impact of $L_{p}$ norm
-For the DDIM model, we train a neural network using the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector v for each image.
-This corresponds to the $L_{1}$ norm.
-Then we change the $L_{p}$ norm of our neural network classifier to see if it has any influence on the results.
-The results are in Figure 12:
-Figure 12: The impact of $L_{p}$ norms. Our algorithm remains effective for $p=1,2,3,4$.
-From the results, we know that our algorithm is effective for $p=1,2,3,4$. We should not choose very large $p$ to train the neural network.
-Impact of Sampling Intervals
+The Impact of Sampling Intervals
 We change the sampling intervals to see if there is any influence on the results.
-In addition to the AUC figures in the main paper, the ASR figures are also plotted in Figure 13.
-From the results, we observe that the different sampling intervals have minimal impact on ASR.
-Figure 13: The impact of sampling intervals. Left: DDIM model on CIFAR-10. Right: Stable Diffusion model on LAION-5. We find that adjusting the sampling interval does not significantly affect the ASR.
+In addition to the AUC figures in the main paper, the ASR figures of DDIM and Diffusion Transformer are also plotted in Figure 13.
+We also plot the AUC and ASR for Stable Diffusion in Figure 14.
+From the results, we observe that our algorithm consistently performs well across different sampling intervals.
+Figure 13: The impact of sampling intervals.
+Left: DDIM model on CIFAR-10.
+Right: Diffusion Transformer on Imagenet.
+We find that adjusting the sampling interval does not significantly affect of the accuracy.
+Figure 14: The impact of sampling intervals on Stable Diffusion. We plot the AUC and ASR metrics, observe that different sampling intervals have minimal impact on it.
+The ROC curves
+Besides the ROC curves in the main paper, we also plot ROC curves for the Diffusion Transformer train on ImageNet $128\times 128$ and the Stable Diffusion train on Laion5 in Figure 15.
+The curves further demonstrate the effectiveness of our method.
+Figure 15: The ROC curves of various setups.
+Left: Diffusion Transformer on ImageNet $128\times 128$.
+Right: The Stable Diffusion on Laion5.
+The curves show that our algorithm outperforms the baseline algorithms.
 Appendix C Proof
@@
 $=\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}(\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t))\,.$
-If we denote the random variable $X_{i}$ represents $\|\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t))\|$, then we consider $S_{n}=\sum_{i=1}^{n}X_{i}$.
-From the assumption we know that all the $X_{i}$ are centered random variable $X$ with finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$.
+If we denote the random variable $X^{i}=(X^{i}_{1},X^{i}_{2},\ldots,X^{i}_{d})$ represents $\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t)$, then we consider $S^{i}_{n}=\sum_{j=1}^{n}X^{i}_{j}$.
+From the assumption we know that all the $X^{i}_{j}$ are random variables with zero expectation and finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$.
 Using the theorem of Chernoff-Cramer method for sums of IID RV [8], we get the following probability inequality for any $\beta>0$:
-$\mathbb{P}(S_{n}\geq\beta)\leq\exp(-n\Psi^{*}_{X}(\frac{\beta}{n}))\,.$
-Therefore, taking the definition of $\|\hat{x}-x\|=\|\frac{1}{n}\sum_{i=1}^{n}(x^{i}-x)\|\leq\frac{\sqrt{1-\bar{\alpha}_{t}}}{n\sqrt{\bar{\alpha}_{t}}}S_{n}$, we get the following bound of the reconstruction error:
-$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq\exp(-n\Psi^{*}_{X}(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}))\,,$
+$\mathbb{P}(|S^{i}_{n}|\geq\beta)\leq\exp(-n\Psi^{*}_{X_{i}}(\frac{\beta}{n}))\,,$
 where $\beta>0$ and $\Psi^{*}_{X}(y)=\sup_{s>0}(sy-\Psi_{X}(s))$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X}$.
+Therefore, denote $S_{n}=(S^{1}_{n},S^{2}_{n},\ldots,S^{d}_{n})$, taking the definition of $\|\hat{x}-x\|=\|\frac{1}{n}\sum_{i=1}^{n}(x^{i}-x)\|=\frac{\sqrt{1-\bar{\alpha}_{t}}}{n\sqrt{\bar{\alpha}_{t}}}\|S_{n}\|$, we get the following bound of the reconstruction error:
+$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq\sum^{d}_{i=1}P(|S^{i}_{n}|\geq\frac{n\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{d(1-\bar{\alpha}_{t}})})\leq d\exp(-n\min_{i}\Psi^{*}_{X_{i}}(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{d(1-\bar{\alpha}_{t}})}))\,,$
 So averaging the randomly reconstructed data from a training set will result in a smaller reconstruction error with high probability of $\Theta(1-\exp(-n))$ when we use a large $n$.
+Appendix D More Results of Variation Images
+In this section, we provide more results of the variation of member and nonmember image when applying to DALL-E 2’s variation API.
+The experimental results are recorded in Figure 16, from which we can see that the variation in the member inputs after applying the DALL-E 2’s variation API is relatively small than non-member inputs.
+Figure 16: More results of the variation images. From the figures we can see that DALL-E 2’s variation API makes less changes to images in member set than non-member set.

### v3

@@
 This makes our algorithm applicable to more setups.
+6 An Application to DALL-E’s API
+In this section, we conduct a small experiment with online API services to test the effectiveness of our algorithm.
+We test with the DALL-E 2 [35] model since DALL-E 2 provides a variation API service.
+We select different thresholds and classify an image with a variation error below the threshold as members and those above the threshold as non-members.
+We then calculate metrics such as AUC and ASR.
+Since the baseline algorithms [30, 6, 20] require intermediate results, we are unable to test these algorithms under the online API setup.
+One challenge with MIA for DALL-E 2 is that it does not disclose its training set.
+However, since it is adept at generating various famous artworks, we select 30 famous paintings from five famous artists: Van Gogh, Monet, Da Vinci, Dali, and Rembrandt, to form our member set.
+We believe that it is reasonable to hypothesize that these artworks are used in DALL-E 2’s training set.
+For constructing the non-member set, we used Stable Diffusion 3 [9] to generate images based on the titles of each painting in the member set.
+The benefit of constructing non-members in this way is that it allows for control over the content of the artwork descriptions, reducing bias caused by content shift.
+Moreover, these generated images are certainly not in the DALL-E 2’s training set.
+Other results of changes to the member and non-member inputs after applying the variation API can be found in Appendix D.
+The results, presented in Table 4, demonstrate that our algorithm achieves a relatively high accuracy under this evaluation method.
+Our observation is illustrated in Figure 17.
+As seen in the figure, for Monet’s iconic painting "Water Lily Pond", the original artwork shows minimal changes when using DALL-E 2’s variation API, retaining most of its main features.
+In contrast, the artwork generated by Stable Diffusion 3 undergoes significant changes, with variations in both the number of flowers and lily pads.
+Therefore, we hypothesize that artworks with smaller changes after API usage are more likely to have appeared in the model’s training set.
+Table 4: The results of applying our algorithm to DALL-E 2’s variation API.
+We assume some famous paintings as members and use Stable Diffusion 3 along with the titles of these artworks to generate corresponding non-members.
+Our algorithm also achieves high accuracy under this setup.
+Since the baseline algorithm requires noise prediction results, we are unable to evaluate it in this black-box setup.
+Metrics $L_{1}$ distance $L_{2}$ distance
+AUC 76.2 88.3
+ASR 74.5 81.4
+Figure 7: Main observation of our attack. DALL-E 2’s variation API makes minimal changes to famous artworks, while nonmember images with similar content undergo significant alterations.
+The results indicate that we can apply our algorithm with online API services.
+We acknowledge that this part of the experimental design has certain limitations.
+Not every famous painting we selected may be present in DALL-E 2’s training set, and our construction of non-members may exhibit some distribution shift relative to the member dataset.
+Here, we aim to provide a real-world application for black-box evaluation, leaving a more comprehensive experimental design as future work.
 7 Conclusion, Limitations and Future Directions

### v4

@@
 Our experimental results consistently outperform previous methods.
+Machine Learning, ICML
 1 Introduction
-Recently, there has been a surge in the popularity of generative models, with diffusion models in particular, gaining huge attention within the AI community [42, 45, 46].
-These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation [14, 44], text-to-image generation [36, 53, 31] and image-to-image generation [37].
-This surge has given rise to powerful AI art models such as DALL-E 2 [35], Stable Diffusion [36], and Imagen [38].
+Recently, there has been a surge in the popularity of generative models, with diffusion models in particular, gaining huge attention within the AI community (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Song et al., 2020b).
+These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation (Ho et al., 2020; Song et al., 2020a), text-to-image generation (Rombach et al., 2022; Yu et al., 2022; Nichol et al., 2021) and image-to-image generation (Saharia et al., 2022a).
+This surge has given rise to powerful AI art models such as DALL-E 2 (Ramesh et al., 2022), Stable Diffusion (Rombach et al., 2022), and Imagen (Saharia et al., 2022b).
 AI-generated art holds a promising future and is expected to have a widespread impact.
@@
 As these models are increasingly deployed across industries, detecting whether a specific piece of content was used in training can help prevent unauthorized use of artistic works, protecting creators’ copyrights and ownership rights.
-This is a classic problem in the field of machine learning, first introduced by [41] and named “membership inference attack”.
+This is a classic problem in the field of machine learning, first introduced by Shokri et al.
+(2017) and named “membership inference attack”.
 A series of studies have been conducted on membership inference attacks against diffusion models.
-[16] was the first to examine this issue, utilizing the loss function values of diffusion models to determine whether an image is in the training set.
-[6]and [20] extended this work by relaxing the assumptions about model access requirements.
+Hu & Pang (2023) was the first to examine this issue, utilizing the loss function values of diffusion models to determine whether an image is in the training set.
+Duan et al.
+(2023) and Kong et al.
+(2023) extended this work by relaxing the assumptions about model access requirements.
 Despite great progress, previous methods are not yet ready for MIA in proprietary diffusion models.
@@
   2. 2.
-    We evaluate our method using DDIM [44] and Stable Diffusion [36] models on classical datasets, including CIFAR10/100 [22], STL10-Unlabeled [2], LAION-5B [40], etc.
+    We evaluate our method using DDIM (Song et al., 2020a) and Stable Diffusion (Rombach et al., 2022) models on classical datasets, including CIFAR10/100 (Krizhevsky et al., 2009), STL10-Unlabeled (Coates et al., 2011), LAION-5B (Schuhmann et al., 2022), etc.
 Our method outperforms the previous methods.
   3. 3.
-    We extend both existing algorithms and our own algorithm to the Diffusion Transformer [33] architecture, implementing the membership inference attack within this model framework for the first time.
+    We extend both existing algorithms and our own algorithm to the Diffusion Transformer (Peebles & Xie, 2023) architecture, implementing the membership inference attack within this model framework for the first time.
 Experimental results demonstrate that our algorithm is consistently effective.
@@
 Diffusion Model
-The diffusion model, initially proposed by [42], has achieved remarkable results in producing high-quality samples across a variety of domains.
-This ranges from image generation [45, 46, 4], audio synthesis [34, 21, 17], and video generation [13, 15, 50], etc.
-Among existing diffusion models, the Denoising Diffusion Probabilistic Model (DDPM) [14] is one of the most frequently adopted.
+The diffusion model, initially proposed by Sohl-Dickstein et al.
+(2015), has achieved remarkable results in producing high-quality samples across a variety of domains.
+This ranges from image generation (Song & Ermon, 2019; Song et al., 2020b; Dhariwal & Nichol, 2021), audio synthesis (Popov et al., 2021; Kong et al., 2020; Huang et al., 2022), and video generation (Ho et al., 2022a, b; Wu et al., 2023), etc.
+Among existing diffusion models, the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) is one of the most frequently adopted.
 This approach introduces a dual-phase process for image generation: initially, a forward process gradually transforms training data into pure noise, followed by a reverse process that meticulously reconstructs the original data from this noise.
-Building on this model, there have been numerous follow-up studies, such as Stable Diffusion [36], which compresses images into a latent space and generates images based on text, and the Denoising Diffusion Implicit Models (DDIM) [44], which removes Gaussian randomness to accelerate the sampling generation process.
+Building on this model, there have been numerous follow-up studies, such as Stable Diffusion (Rombach et al., 2022), which compresses images into a latent space and generates images based on text, and the Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a), which removes Gaussian randomness to accelerate the sampling generation process.
 These advancements demonstrate the versatility and potential of diffusion models.
@@
 The training of diffusion models may involve sensitive datasets like artists’ artworks, which are protected by copyright laws.
-Membership inference attacks, initially introduced by [41], serve as an effective means to detect potential misuse of data without proper authorization.
+Membership inference attacks, initially introduced by Shokri et al.
+(2017), serve as an effective means to detect potential misuse of data without proper authorization.
 Its objective is to ascertain whether a particular data sample participated in the training phase of a target model.
 This approach is instrumental in probing privacy breaches and identifying illicit data utilization.
-Researchers primarily focus on membership inference attacks for classification models [39, 52, 26, 24], embedding models [43, 7, 28], and generative models [10, 12, 1].
-In the domain of membership inference attacks against diffusion models, [51, 16] use a white-box approach, which assumes access to the entire diffusion model and utilizes loss and likelihood to determine whether a sample is in the training set.
-[6, 20, 47] have relaxed these requirements, eliminating the need for the entire model.
+Researchers primarily focus on membership inference attacks for classification models (Salem et al., 2018; Yeom et al., 2018; Long et al., 2018; Li & Zhang, 2021), embedding models (Song & Raghunathan, 2020; Duddu et al., 2020; Mahloujifar et al., 2021), and generative models (Hayes et al., 2017; Hilprecht et al., 2019; Chen et al., 2020).
+In the domain of membership inference attacks against diffusion models, Wu et al.
+(2022); Hu & Pang (2023) use a white-box approach, which assumes access to the entire diffusion model and utilizes loss and likelihood to determine whether a sample is in the training set.
+Duan et al.
+(2023); Kong et al.
+(2023); Tang et al.
+(2023) have relaxed these requirements, eliminating the need for the entire model.
 They leverage the insight that samples within the training set yield more accurate noise predictions, thereby achieving high accuracy in membership inference attack tasks.
 However, they also require the outputs of the U-net, as it is necessary to obtain the noise predictions of intermediate steps.
-Recently, [32] proposed a black-box membership inference attack method against diffusion models.
+Recently, Pang & Wang (2023) proposed a black-box membership inference attack method against diffusion models.
 Their method identifies whether a specific image is in a finetuning dataset of 100 images by calculating the difference between the generated image and the target image, using the corresponding prompt as input.
@@
 In this section, we begin by introducing the notations used for several popular diffusion models.
-We first introduce the Denoising Diffusion Probabilistic Model (DDPM) [14].
-Then, we extend to the Denoising Diffusion Implicit Model (DDIM) [44] and Stable Diffusion [36], which are variants of DDPM used to accelerate image generation or generate images grounded in text descriptions.
-Lastly, we discuss Diffusion Transformer [33], a model that replaces the U-net architecture with a transformer and achieves higher-quality image generation.
+We first introduce the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020).
+Then, we extend to the Denoising Diffusion Implicit Model (DDIM) (Song et al., 2020a) and Stable Diffusion (Rombach et al., 2022), which are variants of DDPM used to accelerate image generation or generate images grounded in text descriptions.
+Lastly, we discuss Diffusion Transformer (Peebles & Xie, 2023), a model that replaces the U-net architecture with a transformer and achieves higher-quality image generation.
 Denoising Diffusion Probabilistic Model (DDPM)
@@
 Unlike DDPM, which requires a large number of denoising steps, DDIM uses a non-Markovian process to accelerate sampling.
-$x_{t-1}=\phi_{\theta}(x_{t},t)=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x_{t},t)\,,$ (1)
+$x_{t-1}=\phi_{\theta}(x_{t},t)$ $=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)$
+$+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x_{t},t)\,,$
 where $\bar{\alpha}_{t}=\prod_{k=0}^{t}\alpha_{k}$, $\alpha_{t}+\beta_{t}=1$ and $\epsilon_{\theta}(x_{t},t)$ is the noise predicted by the model at step $t$.
@@
 Stable Diffusion
-Stable Diffusion leverages a variational autoencoder (VAE) [19] to encode images into a latent space and perform diffusion in this compressed space.
+Stable Diffusion leverages a variational autoencoder (VAE) (Kingma & Welling, 2013) to encode images into a latent space and perform diffusion in this compressed space.
 The model uses a text encoder to guide the diffusion process, enabling text-to-image generation:
@@
 Diffusion Transformer
-Diffusion Transformer leverages the Vision Transformer [5] structure to replace the U-net architecture traditionally used in diffusion models for noise prediction.
+Diffusion Transformer leverages the Vision Transformer (Dosovitskiy, 2020) structure to replace the U-net architecture traditionally used in diffusion models for noise prediction.
 Its training and sampling methods remain consistent with DDIM, with the only difference being the replacement of noise prediction network $\epsilon_{\theta}$ with $\epsilon_{\tilde{\theta}}$, where $\tilde{\theta}$ represents a Vision Transformer-based architecture.
@@
 Most previous works on membership inference attacks against diffusion models aim to prevent data leakage and hence rely on thresholding the model’s training loss.
-For instance, [16] involves a direct comparison of image losses, while [6, 20] evaluates the accuracy of the model’s noise prediction at initial or intermediate steps.
+For instance, Hu & Pang (2023) involves a direct comparison of image losses, while Duan et al.
+(2023); Kong et al.
+(2023) evaluates the accuracy of the model’s noise prediction at initial or intermediate steps.
 However, the required access to the model’s internal U-net structure prevents applications from copyright protection because most servers typically provide only black-box API access.
@@
 Below, we formalize the definition of the variation API used in our algorithm.
-Definition 1 (The variation API).
+Definition 4.1 (The variation API).
 We define the variation API $V_{\theta}(x,t)$ of a model as follows.
 Suppose we have an input image $x$, and the diffusion step of the API is $t$.
-The variation API randomly adds $t$-step Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to the image and denoises it using the DDIM sampling process $\phi_{\theta}(x_{t},t)$ as described in Equation 1, returning the reconstructed image $V_{\theta}(x,t)=\hat{x}$.
+The variation API randomly adds $t$-step Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to the image and denoises it using the DDIM sampling process $\phi_{\theta}(x_{t},t)$, returning the reconstructed image $V_{\theta}(x,t)$.
 The details are as follows:
-$x_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\quad\hat{x}=\Phi_{\theta}(x_{t},0)=\phi_{\theta}(\cdots\phi_{\theta}(\phi_{\theta}(x_{t},t),t-1),0)\,.$ (2)
-This definition aligns with the image-to-image generation method of the diffusion model, making access to this API practical in many setups [27, 37, 49].
+$x_{t}$ $=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,$
+$V_{\theta}(x,t)$ $=\Phi_{\theta}(x_{t},0)=\phi_{\theta}(\cdots\phi_{\theta}(\phi_{\theta}(x_{t},t),t-1),0)\,.$
+This definition aligns with the image-to-image generation method of the diffusion model, making access to this API practical in many setups (Lugmayr et al., 2022; Saharia et al., 2022a; Wu & De la Torre, 2023).
 Some APIs provide the user with a choice of $t$, while others do not and use the default parameter.
 We will discuss the influence of different diffusion steps in Section 5.5, showing that the attack performances are relatively stable and not sensitive to the selection of $t$.
-We also note that for the target model, we can substitute $\phi_{\theta}(x_{t},0)$ with other sampling methods, such as the Euler-Maruyama Method [29] or Variational Diffusion Models [18].
+We also note that for the target model, we can substitute $\phi_{\theta}(x_{t},0)$ with other sampling methods, such as the Euler-Maruyama Method (Mao, 2015) or Variational Diffusion Models (Kingma et al., 2021).
 Figure 1: The overview of ReDiffuse.
@@
 $L(\theta)=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right\|^{2}\right]\,.$
-Denote $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, we assume that the denoise model is expressive enough such that for the input $x_{0}\in\mathbb{R}^{d}$ and time step $t\in\mathcal{T}$, the Jacobian matrix $\nabla_{\theta}\epsilon_{\theta}(x_{t},t)$ is full rank $(\geq d)$.
+Denote $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, we assume that the denoise model is expressive enough(the neural network’s dimensionality significantly exceeds that of the image data) such that for the input $x_{0}\in\mathbb{R}^{d}$ and time step $t\in\mathcal{T}$, the Jacobian matrix $\nabla_{\theta}\epsilon_{\theta}(x_{t},t)$ is full rank $(\geq d)$.
 This suggest that the model can adjust the predicted noise $\epsilon_{\theta}(x_{t},t)$ locally in any direction.
-Then for a well trained model, we would have
-$\nabla_{\theta}L(\theta)=0$ $\implies\nabla_{\theta}\epsilon_{\theta}(x_{t},t)^{T}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right]=0\,,$
-$\implies\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right]=0\,.$
-Intuitively, this is because if the noise prediction from the neural network exhibited high bias, the network could adjust to fit the bias term, further reducing the training loss.
+Then for a well trained model, we would have $\nabla_{\theta}L(\theta)=0$.
+This implies
+$\implies\nabla_{\theta}\epsilon_{\theta}(x_{t},t)^{T}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(x_{t},t\right)\right]=0\,,$
+$\implies\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(x_{t},t\right)\right]=0\,.$
+This is intuitive because if the noise prediction from the neural network exhibits high bias, the network can adjust to fit the bias term, further reducing the training loss.
+Therefore, for images in the training set, we expect the network to provide an unbiased noise prediction.
+Since the noise prediction is typically inaccessible in practical applications, we use the reconstructed sample $\hat{x}$ as a proxy.
+By leveraging the unbiasedness of noise prediction, we show that averaging over multiple independent reconstructed samples $\hat{x}_{i}$ can significantly reduce estimation error (see Theorem 4.2).
+However, for images not in the training set, the neural network may not provide an unbiased prediction at these points.
+The intuition is illustrated in Figure 2.
 Figure 2: The intuition of our algorithm design.
@@
 For member image $x$, the difference $v=x-\hat{x}$ will be smaller after averaging due to $x_{i}$ being an unbiased estimator.
-Therefore, for images in the training set, we expect the network to provide an unbiased noise prediction.
-Since the noise prediction is typically inaccessible in practical applications, we use the reconstructed sample $\hat{x}$ as a proxy.
-Leveraging the unbiasedness of noise prediction, we demonstrate that averaging over multiple independent reconstructed samples $\hat{x}_{i}$ significantly reduces estimation error (see Theorem 1).
-On the other hand, for images that are not in the training set, the neural network may not provide an unbiased prediction at these points.
-We illustrate the intuition in Figure 2.
+Table 1: Comparison of different methods on four datasets for the DDIM model.
+Previous methods require access to the U-Net, whereas our methods do not.
+We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
+Method CIFAR10 CIFAR100 STL10
+Algorithm U-Net AUC ASR TP AUC ASR TP AUC ASR TP
+Loss (Matsumoto et al., 2023) $\Box$ 0.88 0.82 14.2 0.92 0.84 20.9 0.89 0.82 15.6
+SecMI (Duan et al., 2023) $\Box$ 0.95 0.90 40.7 0.96 0.90 44.9 0.94 0.88 26.9
+PIA (Kong et al., 2023) $\Box$ 0.95 0.89 48.7 0.96 0.90 47.0 0.94 0.87 29.8
+PIAN (Kong et al., 2023) $\Box$ 0.95 0.89 50.4 0.91 0.85 39.2 0.92 0.86 28.5
+ReDiffuse $\blacksquare$ 0.96 0.91 40.7 0.98 0.93 48.2 0.96 0.90 31.9
+$\Box$: Require the access of U-Net.  $\blacksquare$: Do not require the access of U-Net.
 With the above intuition, we introduce the details of our algorithm.
@@
 Our algorithm classifies a sample as being in the training set if $D(x,\hat{x})$ is smaller than a threshold $\tau$, where $D(x,\hat{x})$ represents the difference between the two images.
-It can be calculated using traditional functions, such as the SSIM metric [48].
+It can be calculated using traditional functions, such as the SSIM metric (Wang et al., 2004).
 Alternatively, we can train a neural network as a proxy.
@@
 Our algorithm is outlined in Algorithm 1, and we name it ReDiffuse.
-The key ideas of our algorithm are illustrated in Figure 1, and we also provide some theoretical analysis in Theorem 1 to support it.
+The key ideas of our algorithm are illustrated in Figure 1, and we also provide some theoretical analysis in Theorem 4.2 to support it.
 Algorithm 1 MIA through ReDiffuse
@@
   for $k=1,\ldots,n$ do
-     Use the variation API $V_{\theta}$ to generate the variation image $\hat{x}_{k}=V_{\theta}(x,t)$ according to Equation (2).
+     Use the variation API $V_{\theta}$ to generate the variation image $\hat{x}_{k}=V_{\theta}(x,t)$ according to Definition 4.1.
   end for
   Average the reconstructed images from each iteration $\hat{x}=\frac{1}{n}(\hat{x}_{1}+\hat{x}_{2}+\ldots+\hat{x}_{n})$.
-  return  "YES" if the distance between the two images $D(x,\hat{x})$ is less than $\tau$, otherwise "NO".
+  return ”YES” if the distance between the two images $D(x,\hat{x})$ is less than $\tau$, otherwise ”NO”.
 Analysis
@@
 We have the following theorem showing that if we use the variation API to input a member $x\sim D_{\text{training}}$, then the error $\|\hat{x}-x\|$ from our method will be small with high probability.
-Theorem 1.
-Suppose the DDIM model can learn a parameter $\theta$ such that, for any $x\sim D_{\text{training}}$ with dimension $d$, the prediction error $\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)$ is a random variable $X=(X_{1},X_{2},\ldots,X_{d})$ with zero expectation and finite cumulant-generating function for each coordinate [8].
+Theorem 4.2.
+Suppose the DDIM model can learn a parameter $\theta$ such that, for any $x\sim D_{\text{training}}$ with dimension $d$, the prediction error $\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)$ is a random variable $X=(X_{1},X_{2},\ldots,X_{d})$ with zero expectation and finite cumulant-generating function for each coordinate (Durret, 2010).
 Suppose the sampling interval $k$ is equal to the variation API diffusion step $t$.
@@
 We defer the proof of this theorem to Appendix C.
-We note that the unbiased assumption on predicted noise can be strong in practice.
-For experiments, we expect the mean $\|\mu\|$ of predicted noise for trained data is smaller than the mean $\|\mu^{\prime}\|$ from unseen data.
-As a result, the empirical best choice of $n$ would be determined by the gap $\|\mu^{\prime}\|-\|\mu\|$.
 4.3 MIA on Other Diffusion Models
@@
 Again we denote $\hat{x}=V_{\theta}(x,t)$, and the details are as follows:
-$z=\text{Encoder}(x)\,,\quad z_{t}=\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\quad\hat{z}=\Phi_{\theta}(z_{t},0)\,,\quad\hat{x}=\text{Decoder}(\hat{z})\,.$ (3)
+$z$ $=\text{Encoder}(x)\,,\quad z_{t}=\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,$
+$\hat{z}$ $=\Phi_{\theta}(z_{t},0)\,,\quad\hat{x}=\text{Decoder}(\hat{z})\,.$
 This definition aligns with the image generation process of the stable diffusion model.
@@
 5 Experiments
+Table 2: Comparison of different methods for Diffusion Transformer using the same set of metrics as Table 1.
+Previous methods require access to the Vision Transformer, whereas our methods do not.
+We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
+Method ImageNet 128 $\times$ 128 ImageNet 256 $\times$ 256
+Algorithm Transformer AUC ASR TP AUC ASR TP
+Loss (Matsumoto et al., 2023) $\Box$ 0.83 0.76 10.7 0.78 0.70 7.3
+SecMI (Duan et al., 2023) $\Box$ 0.80 0.73 8.3 0.88 0.80 16.3
+PIA (Kong et al., 2023) $\Box$ 0.97 0.92 32.1 0.91 0.85 6.8
+PIAN (Kong et al., 2023) $\Box$ 0.66 0.64 6.2 0.67 0.66 12.8
+ReDiffuse $\blacksquare$ 0.98 0.95 44.1 0.97 0.94 47.3
+$\Box$: Require the access of Transformer.  $\blacksquare$: Do not require the access of Transformer.
+Table 3: Comparison of different methods for Stable Diffusion using the same set of metrics as Table 1.
+Again, previous methods require access to U-Net, whereas our methods do not.
+We use AUC, ASR and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
+Method Laion5 Laion5 with BLIP
+Algorithm U-Net AUC ASR TP AUC ASR TP
+Loss (Matsumoto et al., 2023) $\Box$ 0.62 0.61 13.2 0.62 0.62 13.3
+SecMI (Duan et al., 2023) $\Box$ 0.70 0.65 19.2 0.71 0.66 19.8
+PIA (Kong et al., 2023) $\Box$ 0.70 0.66 19.7 0.73 0.68 20.2
+PIAN (Kong et al., 2023) $\Box$ 0.56 0.53 4.8 0.55 0.51 4.4
+ReDiffuse $\blacksquare$ 0.81 0.75 20.6 0.82 0.75 21.7
+$\Box$: Require the access of U-Net.  $\blacksquare$: Do not require the access of U-Net.
 In this section, we evaluate the performance of our methods across various datasets and settings.
-We follow the same experiment setup in previous papers [6, 20].
+We follow the same experiment setup in previous papers (Duan et al., 2023; Kong et al., 2023).
 The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A.
 5.1 Evaluation Metrics
-We follow the metrics used in previous papers [6, 20], including Area Under Receiver Operating Characteristic (AUC), Attack Success Rate (ASR), the True Positive Rate (TP) when the False Positive Rate is 1%.
+We follow the metrics used in previous papers (Duan et al., 2023; Kong et al., 2023), including Area Under Receiver Operating Characteristic (AUC), Attack Success Rate (ASR), the True Positive Rate (TP) when the False Positive Rate is 1%.
 We also plot the ROC curves.
 5.2 MIA with DDIM Models
-We follow the experimental setup of [6, 20].
-We train a DDIM model on the CIFAR-10/100 [22] and STL10-Unlabeled datasets [2], using the image generation step $T=1000$ and sampling interval $k=100$.
+We follow the experimental setup of (Duan et al., 2023; Kong et al., 2023).
+We train a DDIM model on the CIFAR-10/100 (Krizhevsky et al., 2009) and STL10-Unlabeled datasets (Coates et al., 2011), using the image generation step $T=1000$ and sampling interval $k=100$.
 For all the datasets, we randomly select 50% of the training samples to train the model and denote them as members.
 The remaining 50% are utilized as nonmembers.
-We use [30], [6], [20] as our baseline methods.
+We use (Matsumoto et al., 2023), (Duan et al., 2023), (Kong et al., 2023) as our baseline methods.
 We fix the diffusion step at $t=200$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
 We will discuss the impact of the diffusion step and the average number in Section 5.5.
-Table 1: Comparison of different methods on four datasets for the DDIM model. We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
-Method CIFAR10 CIFAR100 STL10
-Algorithm U-Net AUC ASR TP AUC ASR TP AUC ASR TP
-Loss [30] $\Box$ 0.88 0.82 14.2 0.92 0.84 20.9 0.89 0.82 15.6
-SecMI [6] $\Box$ 0.95 0.90 40.7 0.96 0.90 44.9 0.94 0.88 26.9
-PIA [20] $\Box$ 0.95 0.89 48.7 0.96 0.90 47.0 0.94 0.87 29.8
-PIAN [20] $\Box$ 0.95 0.89 50.4 0.91 0.85 39.2 0.92 0.86 28.5
-ReDiffuse $\blacksquare$ 0.96 0.91 40.7 0.98 0.93 48.2 0.96 0.90 31.9
-$\Box$: Require the access of U-Net.  $\blacksquare$: Do not require the access of U-Net.
-For the difference function $D(x,\hat{x})$, following the setup in [6], we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image.
-Using the ResNet-18 network [11] and denoting it as $f_{R}$, we perform binary classification on these difference vectors.
+For the difference function $D(x,\hat{x})$, following the setup in (Duan et al., 2023), we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image.
+Using the ResNet-18 network (He et al., 2016) and denoting it as $f_{R}$, we perform binary classification on these difference vectors.
 We use 20% of the data as the training set and obtained the label of each difference vector being classified as a member or nonmember.
@@
 5.3 MIA with Diffusion Transformers
-We train a diffusion transformer model on the ImageNet [3] dataset following the setup of [33].
+We train a diffusion transformer model on the ImageNet (Deng et al., 2009) dataset following the setup of (Peebles & Xie, 2023).
 We randomly select 100,000 images from the ImageNet training set to train the model with resolutions of either $128\times 128$ or $256\times 256$.
@@
 We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
-We use [30], [6], [20] as our baseline methods.
+We use (Matsumoto et al., 2023), (Duan et al., 2023), (Kong et al., 2023) as our baseline methods.
 Since these work did not study the case of Diffusion Transformers, we integrate their algorithms into the DiT framework for evaluation.
-For the difference function $D(x,\hat{x})$, following the setup in [6, 20], we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
-Table 2: Comparison of different methods for Diffusion Transformer using the same set of metrics as Table 1.
-Previous methods require access to the Vision Transformer, whereas our methods do not.
-We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
-Method ImageNet 128 $\times$ 128 ImageNet 256 $\times$ 256
-Algorithm Transformer AUC ASR TP AUC ASR TP
-Loss [30] $\Box$ 0.83 0.76 10.7 0.78 0.70 7.3
-SecMI [6] $\Box$ 0.80 0.73 8.3 0.88 0.80 16.3
-PIA [20] $\Box$ 0.97 0.92 32.1 0.91 0.85 6.8
-PIAN [20] $\Box$ 0.66 0.64 6.2 0.67 0.66 12.8
-ReDiffuse $\blacksquare$ 0.98 0.95 44.1 0.97 0.94 47.3
-$\Box$: Require the access of Transformer.  $\blacksquare$: Do not require the access of Transformer.
-We also plot ROC curves for the DDIM train on CIFAR-10 and the Diffusion Transformer train on ImageNet $256\times 256$ in Figure 3. The curves further demonstrate the effectiveness of our method.
-Figure 3: The ROC curves of various setups.
-Left: DDIM model on CIFAR-10.
-Right: Diffusion Transformer on ImageNet $256\times 256$.
-The curves show that our algorithm outperforms the baseline algorithms.
+For the difference function $D(x,\hat{x})$, following the setup in (Duan et al., 2023; Kong et al., 2023), we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
+We also plot ROC curves for the DDIM train on CIFAR-10 and the Diffusion Transformer train on ImageNet in Appendix B. The curves further demonstrate the effectiveness of our method.
 5.4 MIA with the Stable Diffusion Model
 We conduct experiments on the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications.
-We follow the experiment setup of [6, 20], use the LAION-5B dataset [40] as member and COCO2017-val [25] as non-member.
+We follow the experiment setup of (Duan et al., 2023; Kong et al., 2023), use the LAION-5B dataset (Schuhmann et al., 2022) as member and COCO2017-val (Lin et al., 2014) as non-member.
 We randomly select 2500 images in each dataset.
-We test two scenarios: Knowing the ground truth text, which we denote as Laion5; Not knowing the ground truth text and generating text through BLIP [23], which we denote as Laion5 with BLIP.
-For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric [48] to measure the differences between two images.
+We test two scenarios: Knowing the ground truth text, which we denote as Laion5; Not knowing the ground truth text and generating text through BLIP (Li et al., 2022), which we denote as Laion5 with BLIP.
+For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric (Wang et al., 2004) to measure the differences between two images.
 The results, presented in Table 3, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%.
 Notably, our methods do not require access to U-Net.
-Table 3: Comparison of different methods for Stable Diffusion using the same set of metrics as Table 1.
-Again, previous methods require access to U-Net, whereas our methods do not.
-We use AUC, ASR and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
-Method Laion5 Laion5 with BLIP
-Algorithm U-Net AUC ASR TP AUC ASR TP
-Loss [30] $\Box$ 0.62 0.61 13.2 0.62 0.62 13.3
-SecMI [6] $\Box$ 0.70 0.65 19.2 0.71 0.66 19.8
-PIA [20] $\Box$ 0.70 0.66 19.7 0.73 0.68 20.2
-PIAN [20] $\Box$ 0.56 0.53 4.8 0.55 0.51 4.4
-ReDiffuse $\blacksquare$ 0.81 0.75 20.6 0.82 0.75 21.7
-$\Box$: Require the access of U-Net.  $\blacksquare$: Do not require the access of U-Net.
 5.5 Ablation Studies
@@
 Figure 3: The impact of average numbers.
-Left: DDIM model on CIFAR-10.
-Right: Diffusion Transformer model on Imagenet.
-In both cases, averaging multiple independent samples proves to be effective in further improving the overall performance of our algorithm, which validates the intuition of our algorithm design.
+We train DDIM model on CIFAR-10.
+Averaging multiple independent samples proves to be effective in further improving the overall performance of our algorithm, which validates the intuition of our algorithm design.
 The Impact of Diffusion Steps
@@
 Figure 5: The impact of sampling intervals.
-Left: DDIM model on CIFAR-10.
-Right: Stable Diffusion model on LAION-5.
+We train DDIM model on CIFAR-10.
 We find that adjusting the sampling interval has a relatively small influence in the first case and does not affect our method in the latter case.
@@
 In this section, we conduct a small experiment with online API services to test the effectiveness of our algorithm.
-We test with the DALL-E 2 [35] model since DALL-E 2 provides a variation API service.
+We test with the DALL-E 2 (Ramesh et al., 2022) model since DALL-E 2 provides a variation API service.
 We select different thresholds and classify an image with a variation error below the threshold as members and those above the threshold as non-members.
 We then calculate metrics such as AUC and ASR.
-Since the baseline algorithms [30, 6, 20] require intermediate results, we are unable to test these algorithms under the online API setup.
+Since the baseline algorithms (Matsumoto et al., 2023; Duan et al., 2023; Kong et al., 2023) require intermediate results, we are unable to test these algorithms under the online API setup.
 One challenge with MIA for DALL-E 2 is that it does not disclose its training set.
@@
 We believe that it is reasonable to hypothesize that these artworks are used in DALL-E 2’s training set.
-For constructing the non-member set, we used Stable Diffusion 3 [9] to generate images based on the titles of each painting in the member set.
+For constructing the non-member set, we used Stable Diffusion 3 (Esser et al., 2024) to generate images based on the titles of each painting in the member set.
 The benefit of constructing non-members in this way is that it allows for control over the content of the artwork descriptions, reducing bias caused by content shift.
 Moreover, these generated images are certainly not in the DALL-E 2’s training set.
-Other results of changes to the member and non-member inputs after applying the variation API can be found in Appendix D.
 The results, presented in Table 4, demonstrate that our algorithm achieves a relatively high accuracy under this evaluation method.
 Our observation is illustrated in Figure 6.
-As seen in the figure, for Monet’s iconic painting "Water Lily Pond", the original artwork shows minimal changes when using DALL-E 2’s variation API, retaining most of its main features.
+As seen in the figure, for Monet’s iconic painting ”Water Lily Pond”, the original artwork shows minimal changes when using DALL-E 2’s variation API, retaining most of its main features.
 In contrast, the artwork generated by Stable Diffusion 3 undergoes significant changes, with variations in both the number of flowers and lily pads.
 Therefore, we hypothesize that artworks with smaller changes after API usage are more likely to have appeared in the model’s training set.
+Other results of changes to the member and non-member inputs after applying the variation API can be found in Appendix D.
 Table 4: The results of applying our algorithm to DALL-E 2’s variation API.
@@
 Our algorithm also achieves high accuracy under this setup.
-Since the baseline algorithm requires noise prediction results, we are unable to evaluate it in this black-box setup.
 Metrics $L_{1}$ distance $L_{2}$ distance
@@
 ASR 74.5 81.4
-Figure 7: Main observation of our attack. DALL-E 2’s variation API makes minimal changes to famous artworks, while nonmember images with similar content undergo significant alterations.
+Figure 6: Main observation of our attack.
+DALL-E 2’s variation API makes minimal changes to famous artworks, while nonmember images with similar content undergo significant alterations.
+More examples of changes to the member and non-member inputs after applying the variation API can be found in Appendix D.
 The results indicate that we can apply our algorithm with online API services.
@@
 Improving the interpretability of our method and extending it to other generative models are also valuable directions for further research.
+Impact Statement
+This paper addresses the critical issue of protecting intellectual property and data privacy in the era of rapidly advancing generative models.
+By proposing a novel membership inference attack (MIA) method, our work aims to identify whether specific artworks have been used during the training of diffusion models.
+This has significant implications for safeguarding creators’ copyrights, detecting unauthorized use of artistic works, and promoting ethical practices in the development and deployment of generative models.
+While our method is designed with privacy protection in mind, we acknowledge the potential for misuse.
+We hope this research serves as a foundation for further discussions on data privacy and ethical considerations, encouraging responsible use and development of diffusion models in the community.
 Appendix A Datasets, Models and Hyperparameters
 We use NVIDIA RTX 6000 graphics cards for all our experiments.
-For DDIM, we follow the training hyperparameters of [6] to train a DDIM model on the CIFAR-10/100 [22] and STL10-Unlabeled [2] datasets, using 1000 image generation steps ($T=1000$) and a sampling interval of $k=100$.
+For DDIM, we follow the training hyperparameters of (Duan et al., 2023) to train a DDIM model on the CIFAR-10/100 (Krizhevsky et al., 2009) and STL10-Unlabeled (Coates et al., 2011) datasets, using 1000 image generation steps ($T=1000$) and a sampling interval of $k=100$.
 The training iterations are set to 800,000.
@@
 The remaining 50% of the training samples are used as nonmembers.
-We use [30, 6, 20] as our baseline methods.
+We use (Matsumoto et al., 2023; Duan et al., 2023; Kong et al., 2023) as our baseline methods.
 We use their official code repositories and apply the optimal hyperparameters from their papers.
 For our algorithm, we fix the diffusion step at $t=200$ and independently call the variation API 10 times to average the output images as $\hat{x}$.
-For the Diffusion Transformers, we train the model on the ImageNet [3] dataset, using 1000 image generation steps ($T=1000$), while for other training hyperparameters, we follow the setup of [33].
+For the Diffusion Transformers, we train the model on the ImageNet (Deng et al., 2009) dataset, using 1000 image generation steps ($T=1000$), while for other training hyperparameters, we follow the setup of (Peebles & Xie, 2023).
 We randomly select 100,000 images from the ImageNet training set to train the model at resolutions of either $128\times 128$ or $256\times 256$.
@@
 We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to average the output images as $\hat{x}$.
-We also use [30, 6, 20] as our baseline methods.
+We also use (Matsumoto et al., 2023; Duan et al., 2023; Kong et al., 2023) as our baseline methods.
 Since these papers do not study Diffusion Transformers, we adapt their algorithms to the DiT framework for evaluation.
@@
 For the Stable Diffusion experiments, we use the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications.
-We follow the experimental setup of [6, 20], using an image generation step of $T=1000$ and a sampling interval of $k=10$.
-We use the LAION-5B dataset [40] as the member set and COCO2017-val [25] as the non-member set.
+We follow the experimental setup of (Duan et al., 2023; Kong et al., 2023), using an image generation step of $T=1000$ and a sampling interval of $k=10$.
+We use the LAION-5B dataset (Schuhmann et al., 2022) as the member set and COCO2017-val (Lin et al., 2014) as the non-member set.
 We randomly select 2500 images from each dataset.
-We test two scenarios: knowing the ground truth text, denoted as Laion; and generating text through BLIP [23], denoted as Laion with BLIP.
-We use the hyperparameters from the papers [6, 20] to run the baseline methods.
-For our algorithms ReDiffuse, we fix the diffusion step at $t=10$ to call the variation API and directly use the SSIM metric [48] to measure the differences between two images.
+We test two scenarios: knowing the ground truth text, denoted as Laion; and generating text through BLIP (Li et al., 2022), denoted as Laion with BLIP.
+We use the hyperparameters from the papers (Duan et al., 2023; Kong et al., 2023) to run the baseline methods.
+For our algorithms ReDiffuse, we fix the diffusion step at $t=10$ to call the variation API and directly use the SSIM metric (Wang et al., 2004) to measure the differences between two images.
 Other hyperparameters remain the same as in the baseline methods.
@@
 The ROC curves
-Besides the ROC curves in the main paper, we also plot ROC curves for the Diffusion Transformer train on ImageNet $128\times 128$ and the Stable Diffusion train on Laion5 in Figure 16.
-The curves further demonstrate the effectiveness of our method.
+We plot ROC curves for the DDIM train on CIFAR-10 and Diffusion Transformer train on ImageNet $256\times 256$ in Figure 15. The curves further demonstrate the effectiveness of our method.
 Figure 15: The ROC curves of various setups.
-Left: Diffusion Transformer on ImageNet $128\times 128$.
-Right: The Stable Diffusion on Laion5.
+Left: DDIM model on CIFAR-10.
+Right: Diffusion Transformer on ImageNet $256\times 256$.
 The curves show that our algorithm outperforms the baseline algorithms.
 Appendix C Proof
-In this section, we implement the proof of the theorem. See 1
+In this section, we implement the proof of the theorem. See 4.2
 Proof.
@@
 From the assumption we know that all the $X^{i}_{j}$ are random variables with zero expectation and finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$.
-Using the theorem of Chernoff-Cramer method for sums of IID RV [8], we get the following probability inequality for any $\beta>0$:
+Using the theorem of Chernoff-Cramer method for sums of IID RV (Durret, 2010), we get the following probability inequality for any $\beta>0$:
 $\mathbb{P}(|S^{i}_{n}|\geq\beta)\leq\exp(-n\Psi^{*}_{X_{i}}(\frac{\beta}{n}))\,,$

### v5

@@
 This surge has given rise to powerful AI art models such as DALL-E 2 (Ramesh et al., 2022), Stable Diffusion (Rombach et al., 2022), and Imagen (Saharia et al., 2022b).
-AI-generated art holds a promising future and is expected to have a widespread impact.
+AI-generated art has a promising future and is expected to have widespread impact.
 Effective training of diffusion models requires high-quality data.
@@
 (2023) and Kong et al.
-(2023) extended this work by relaxing the assumptions about model access requirements.
+(2023) extended this work by relaxing assumptions about the access requirements of the model.
 Despite great progress, previous methods are not yet ready for MIA in proprietary diffusion models.
@@
 4.2 Algorithm
-[…]
+In this section, we present the intuition of our algorithm.
+We denote $\|\cdot\|$ as the $L_{2}$ operator norm of a vector and $\mathcal{T}=\{1,2,\ldots,T\}$ as the set of diffusion steps.
+The key insight is derived from the training loss function of a fixed sample $x_{0}$ and a time step $t\in\mathcal{T}$:
 $L(\theta)=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right\|^{2}\right]\,.$
 Denote $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, we assume that the denoise model is expressive enough (the neural network’s dimensionality significantly exceeds that of the image data) such that for the input $x_{0}\in\mathbb{R}^{d}$ and time step $t\in\mathcal{T}$, the Jacobian matrix $\nabla_{\theta}\epsilon_{\theta}(x_{t},t)$ is full rank $(\geq d)$.
-This suggest that the model can adjust the predicted noise $\epsilon_{\theta}(x_{t},t)$ locally in any direction.
+This suggests that the model can adjust the predicted noise $\epsilon_{\theta}(x_{t},t)$ locally in any direction.
 Then for a well trained model, we would have $\nabla_{\theta}L(\theta)=0$.
@@
 $\implies\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(x_{t},t\right)\right]=0\,.$
-This is intuitive because if the noise prediction from the neural network exhibits high bias, the network can adjust to fit the bias term, further reducing the training loss.
+This is intuitive because if the neural network noise prediction exhibits a high bias, the network can adjust to fit the bias term, further reducing the training loss.
 Therefore, for images in the training set, we expect the network to provide an unbiased noise prediction.
@@
 $f(x)=\mathbf{1}\left[D(x,\hat{x})<\tau\right]\,.$
-Our algorithm classifies a sample as being in the training set if $D(x,\hat{x})$ is smaller than a threshold $\tau$, where $D(x,\hat{x})$ represents the difference between the two images.
+Our algorithm classifies a sample as being in the training set if D(x,x^)D($\hat{x}$) is smaller than a threshold $\tau$, where D(x,x^)D($\hat{x}$) represents the difference between the two images.
 It can be calculated using traditional functions, such as the SSIM metric (Wang et al., 2004).
 Alternatively, we can train a neural network as a proxy.
-In Section 5, we will introduce the details of $D(x,\hat{x})$ used in our experiment.
+In Section 5, we will introduce the details of D(x,x^)D($\hat{x}$) used in our experiment.
 Our algorithm is outlined in Algorithm 1, and we name it ReDiffuse.
@@
  Average the reconstructed images from each iteration $\hat{x}=\frac{1}{n}(\hat{x}_{1}+\hat{x}_{2}+\ldots+\hat{x}_{n})$.
-  return ”YES” if the distance between the two images $D(x,\hat{x})$ is less than $\tau$, otherwise ”NO”.
+return ”YES” if the distance between the two images D(x,x^)D($\hat{x}$) is less than $\tau$, otherwise ”NO”.
 Analysis
@@
 where $\beta>0$ and $\Psi^{*}_{X_{i}}$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X_{i}}$.
-The theorem suggests that averaging the randomly variational data from a training set will result in a smaller relative error with high probability when we use a large $n$.
+The theorem suggests that averaging the random variational data from a training set will result in a lower relative error with high probability when we use a large $n$.
 We defer the proof of this theorem to Appendix C.
@@
 $z$ $=\text{Encoder}(x)\,,\quad z_{t}=\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,$,
-$\hat{z}$ $=\Phi_{\theta}(z_{t},0)\,,\quad\hat{x}=\text{Decoder}(\hat{z})\,.$
+[…]
 This definition aligns with the image generation process of the stable diffusion model.
@@
 We follow the same experiment setup in previous papers (Duan et al., 2023; Kong et al., 2023).
-The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A.
+The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A [^1] Code is available at https://github.com/lijingwei0502/diffusion_mia..
 5.1 Evaluation Metrics
@@
 We will discuss the impact of the diffusion step and the average number in Section 5.5.
-For the difference function $D(x,\hat{x})$, following the setup in (Duan et al., 2023), we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image.
+For the difference function D(x,x^)D($\hat{x}$), following the setup in (Duan et al., 2023), we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image.
 Using the ResNet-18 network (He et al., 2016) and denoting it as $f_{R}$, we perform binary classification on these difference vectors.
@@
 For the membership inference attack setup, 1000 images are randomly chosen from our training set as the member set, and another 1000 images are randomly selected from the ImageNet validation set as the non-member set.
-We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
+We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and we independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
 We use (Matsumoto et al., 2023), (Duan et al., 2023), (Kong et al., 2023) as our baseline methods.
 Since these work did not study the case of Diffusion Transformers, we integrate their algorithms into the DiT framework for evaluation.
-For the difference function $D(x,\hat{x})$, following the setup in (Duan et al., 2023; Kong et al., 2023), we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
+For the difference function D(x,x^)D($\hat{x}$), following the setup in (Duan et al., 2023; Kong et al., 2023), we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
 We also plot ROC curves for the DDIM train on CIFAR-10 and the Diffusion Transformer train on ImageNet in Appendix B. The curves further demonstrate the effectiveness of our method.
@@
 We test two scenarios: Knowing the ground truth text, which we denote as Laion5; Not knowing the ground truth text and generating text through BLIP (Li et al., 2022), which we denote as Laion5 with BLIP.
-For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric (Wang et al., 2004) to measure the differences between two images.
+For the difference function D(x,x^)D($\hat{x}$), since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric (Wang et al., 2004) to measure the differences between two images.
 The results, presented in Table 3, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%.
@@
 We randomly select 100,000 images from the ImageNet training set to train the model at resolutions of either $128\times 128$ or $256\times 256$.
-For the $128\times 128$ image size, we use 160,000 training iterations.
+For the $128\times 128$ image size, we use 200,000 training iterations.
 For the $256\times 256$ image size, we use 300,000 training iterations.
@@
 For the membership inference attack setup, 1000 images are randomly selected from our training set as the member set, and another 1000 images are randomly chosen from the ImageNet validation set as the non-member set.
-We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to average the output images as $\hat{x}$.
+We fix the diffusion step at $t=100$ and the DDIM step at $k=50$, and independently call the variation API 10 times to average the output images as $\hat{x}$.
 We also use (Matsumoto et al., 2023; Duan et al., 2023; Kong et al., 2023) as our baseline methods.
@@
 Appendix C Proof
-In this section, we implement the proof of the theorem. See 4.2
+In this section, we present the proof of the theorem. See 4.2
 Proof.
 We denote $x^{i}$ as the $i$-th output image of $x$ using the variation API.
-We denote the $i$-th Gaussian noise as $\epsilon^{i}$ and the forward process incur $x^{i}_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i}$.
+We denote the $i$-th Gaussian noise as $\epsilon^{i}$ and the forward process yields $x^{i}_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i}$.
 As the sampling interval is equal to the variation API diffusion step $t$, we have
@@
 If we denote the random variable $X^{i}=(X^{i}_{1},X^{i}_{2},\ldots,X^{i}_{d})$ represents $\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t)$, then we consider $S^{i}_{n}=\sum_{j=1}^{n}X^{i}_{j}$.
-From the assumption we know that all the $X^{i}_{j}$ are random variables with zero expectation and finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$.
-Using the theorem of Chernoff-Cramer method for sums of IID RV (Durret, 2010), we get the following probability inequality for any $\beta>0$:
+From the assumption we know that all the $X^{i}_{j}$ are i.i.d.
+random variables(with the same distribution of $X_{i}$ in Theorem 4.2) with zero expectation and finite cumulant-generating function $\Psi_{X_{i}}(s)=\log E[e^{sX_{i}}]<+\infty$.
+Using the theorem of Cramér–Chernoff method for sums of i.i.d.
+random variables (Durret, 2010), we get the following probability inequality for any $\beta>0$:
 $\mathbb{P}(|S^{i}_{n}|\geq\beta)\leq\exp(-n\Psi^{*}_{X_{i}}(\frac{\beta}{n}))\,$,
-where $\beta>0$ and $\Psi^{*}_{X}(y)=\sup_{s>0}(sy-\Psi_{X}(s))$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X}$.
+where $\beta>0$ and $\Psi^{*}_{X_{i}}(y)=\sup_{s>0}(sy-\Psi_{X_{i}}(s))$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X_{i}}$.
 Therefore, denote $S_{n}=(S^{1}_{n},S^{2}_{n},\ldots,S^{d}_{n})$, taking the definition of $\|\hat{x}-x\|=\|\frac{1}{n}\sum_{i=1}^{n}(x^{i}-x)\|=\frac{\sqrt{1-\bar{\alpha}_{t}}}{n\sqrt{\bar{\alpha}_{t}}}\|S_{n}\|$, we get the following bound of the reconstruction error:
 $\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq\sum^{d}_{i=1}P(|S^{i}_{n}|\geq\frac{n\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{d(1-\bar{\alpha}_{t}})})\leq d\exp(-n\min_{i}\Psi^{*}_{X_{i}}(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{d(1-\bar{\alpha}_{t}})}))\,$,
-So averaging the randomly reconstructed data from a training set will result in a smaller reconstruction error with high probability of $\Theta(1-\exp(-n))$ when we use a large $n$.
+So averaging the randomly reconstructed data from a training set will result in a smaller reconstruction error with high probability of $1-\mathcal{O}(\exp(-n))$ when we use a large $n$.
 Appendix D More Results of Variation Images